### v3

# VidCRAFT3: Camera, Object, and Lighting Control for Image-to-Video Generation

Abstract
Recent image-to-video generation methods have demonstrated success in enabling control over one or two visual elements, such as camera motion or object motion. However, these methods are unable to offer control over multiple visual elements due to limitations in data and network efficacy. In this paper, we introduce VidCRAFT3, a novel framework for precise image-to-video generation that enables control over camera motion, object motion, and lighting direction simultaneously. VidCRAFT3 integrates three core components: Image2Cloud generates 3D point cloud from a reference image; ObjMotionNet encodes sparse object trajectories using multi-scale optical flow features; and Spatial Triple-Attention Transformer incorporates lighting direction embeddings via parallel cross-attention modules. Additionally, we introduce the VideoLightingDirection dataset, providing synthetic yet realistic video clips with accurate per-frame lighting direction annotations, effectively mitigating the lack of annotated real-world datasets. We further adopt a three-stage training strategy, ensuring robust learning even without joint multi-element annotations. Extensive experiments show that VidCRAFT3 produces high-quality video content, outperforming state-of-the-art methods in control granularity and visual coherence. Code and data will be publicly available.
Figure 1: VidCRAFT3 is the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction. It offers user-friendly control over camera motion (a trajectory in blue), object motion (sparse trajectories in red), and lighting direction. VidCRAFT3 can take any combination of supported control signals and deliver fine-grained and faithful generation results.
1 Introduction
Image-to-video (I2V) generation is a powerful technique that brings still images to life, enabling widespread applications in the fields of content creation, advertising, and art animation. Typically, I2V generation aims to animate a reference image according to user-defined control signals, such as text, object motion, and camera motion, while maintaining the quality and natural dynamics of the generated video. Driven by advances in diffusion-based generative models and extensive web-scale data [2], recent methods have significantly enhanced the ability to generate temporally coherent and visually compelling videos from limited inputs such as a single image or sparse annotations [15, 6, 60, 14, 57, 17, 28, 13]. However, achieving precise and simultaneous control over multiple video attributes, such as camera motion, object motion, and lighting direction, remains a critical yet challenging objective.
Existing approaches typically address these control attributes independently or in a partially integrated manner. Methods focusing exclusively on camera control, such as CameraCtrl [17] and ViewCrafter, do not adequately handle detailed object motion or dynamic lighting variations. Similarly, object-motion-centric frameworks, including DragAnything [58], lack integrated capabilities for precise camera and lighting control. For lighting control, techniques like DILightNet [66] and NeuLighting [27] provide lighting control but often lack generalizability due to reliance on specific categories (e.g., human faces) and object-specific annotations. Other methods [27, 68, 22, 68, 22] also primarily target narrow domains, such as human faces, further restricting their applicability to open-domain scenarios.
Simultaneous and fine-grained control of camera motion, object motion, and lighting direction introduces multiple significant challenges: (1) Accurate camera motion control from single-image inputs requires robust extraction of sufficient 3D geometric information. (2) Realistic and detailed object motion control demands effective representation and refinement of sparse motion trajectories without compromising visual fidelity. (3) Dynamic lighting control necessitates integrating illumination adjustments coherently with both camera and object motion to ensure temporal consistency and visual realism.
To overcome these challenges, we propose VidCRAFT3, which integrates three core components. First, the Image2Cloud module leverages DUSt3R [54] to reconstruct a 3D point cloud from a single reference image, enabling precise camera motion control by rendering the point cloud along user-defined camera motion trajectory. Second, ObjMotionNet encodes sparse object motion trajectories by extracting multi-scale motion features from Gaussian-smoothed optical flow maps to guide realistic object motion. Third, Spatial Triple-Attention Transformer integrates lighting direction control by integrating lighting embedding with image-text embeddings through parallel cross-attention layers, ensuring consistent illumination effects. The capabilities of these modules are shown in Fig. 1, showcasing VidCRAFT3’s control over camera motion, object motion, and lighting direction. To address data scarcity, we introduce the VideoLightingDirection (VLD) dataset, providing synthetic yet highly realistic video sequences with accurate per-frame lighting annotations. Additionally, we develop a three-stage training strategy that enables effective learning without joint multi-element annotations.
The main contributions of this paper are: (1) VidCRAFT3 is the first I2V framework to achieve simultaneous control over camera motion, object motion, and lighting direction through a disentangled architecture combining Image2Cloud, ObjMotionNet, and Spatial Triple-Attention Transformer. (2) VLD dataset provides synthetic videos accompanied by accurate per-frame lighting direction annotations, effectively addressing the scarcity of annotated real-world datasets. (3) A three-stage training strategy enables robust multi-element control without requiring real-world videos annotated with camera, object, and lighting signals. (4) Extensive experiments demonstrate that VidCRAFT3 achieves state-of-the-art performance surpassing existing methods in terms of control precision, visual quality, and generalization capabilities.
2 Related Work
2.1 Image-to-video Generation
Image-to-video (I2V) generation [15, 6, 7, 60, 14, 16] aims to animate static images into dynamic videos while preserving visual content and introducing realistic motion. Recent advances in diffusion models [48, 20] have revolutionized video generation by extending pre-trained Text-to-Image (T2I) models like AnimateDiff [15] to incorporate temporal dimensions for motion generation. These methods integrate the input image as a condition, either through CLIP-based [42] image embeddings or by concatenating the image with noisy latent. For example, VideoCrafter1 [6], DynamiCrafter [60], and I2V-Adapter [14] use dual cross-attention layers to fuse image embeddings with noisy frames, ensuring spatial-aligned guidance. Similarly, Stable Video Diffusion (SVD) [4] replaces text embeddings with CLIP image embeddings, maintaining semantic consistency in an image-only manner. Another line of work, exemplified by SEINE [8], DynamiCrafter [60] and PixelDance [67], expands the input channels of diffusion models to concatenate the static image with noisy latents, effectively injecting image information into the model. However, these methods preserve input image fidelity while generating dynamic videos but often struggle with fine-grained details due to reliance on global conditions.
2.2 Motion-controlled Video Generation
Motion-controlled video generation [57, 17, 64] focuses on creating high-fidelity videos with user-defined motion dynamics. Existing methods can be categorized into camera motion control, object motion control, and joint motion control. For camera motion control [17, 1, 26, 21, 56], CamI2V [72], CameraCtrl [17], and CamCo [61] encode camera parameters into Plücker embeddings for more discriminative motion representation. ViewCrafter [65] and I2VControl-Camera [11] use 3D representations of reference images to render point clouds and generate partial frames. For object motion control [32, 46, 55, 23, 30, 71, 74, 62, 37, 19, 12, 34, 59], existing methods employ diverse strategies to guide object movements in video synthesis. DragNUWA [64], Image Conductor [28], DragAnything [58], and MotionBridge [50] utilize sparse optical flow to control object motion, while MOFA-Video [36] and Motion-I2V [45] predict dense motion fields to control object motion. Alternatively, Boximator [52] and Direct-a-Video [63] rely on bounding box trajectories to guide object motion. LeviTor [51] combines depth and K-means clustering for precise 3D trajectory control. Current research on Joint Motion Control [63, 57, 13, 9], aiming to simultaneously control camera and object motions, remains limited. MotionCtrl [57] designs specialized modules for precise control, while Motion Prompting [13] uses motion tracks for flexible guidance. Perception-as-Control [9] further enhances control with 3D-aware motion representations.
Unlike existing methods, we propose VidCRAFT3, the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction. By combining 3D point cloud rendering, trajectory learning, and Spatial Triple-Attention Transformer, our approach effectively decouples these elements, ensuring temporal consistency and enhanced realism in complex scenes.
2.3 Lighting-controlled Visual Generation
Lighting-controlled visual generation aims to manipulate illumination while preserving scene geometry and materials. Previous methods primarily focus on portrait lighting [49, 73, 43, 38, 35, 25, 44, 47, 18, 40, 70], laying the foundation for effective and accurate illumination modeling. Recent advances in diffusion models have significantly improved the quality and flexibility of lighting control. Methods like DiLightNet [66] and GenLit [3] achieve fine-grained and realistic relighting through radiance hints and SVD, respectively. Facial relighting methods, including DifFRelight [18] and DiFaReli [40], produce high-quality portrait images. Frameworks like NeuLighting [27] focus on outdoor scenes using unconstrained photo collections, while GSR [39] combines diffusion models with neural radiance fields for realistic 3D-aware relighting. IC-Light [69] proposes imposing consistent light transport during training. Extending lighting control to video [70] introduces challenges such as temporal consistency and dynamic lighting effects. Recent techniques leverage 3D-aware generative models for temporally consistent relighting, as seen in EdgeRelight360 [29] and ReliTalk [41]. Neural rendering approaches [5, 68] use datasets like dynamic OLAT for high-quality portrait video relighting, while reflectance field-based methods [22] infer lighting from exemplars. Despite these advancements, most prior works focus on portraits and HDR lighting conditions. In contrast, VidCRAFT3 targets general scenes, enabling interactive adjustment of lighting directions.
3 Method
3.1 Overview
Figure 2: Architecture of VidCRAFT3 for controlled image-to-video generation. The model builds on Video Diffusion Model (VDM) and consists of three main components: the Image2Cloud reconstructs 3D point cloud from a single reference image and generates point cloud renderings along user-defined camera motion trajectory; the ObjMotionNet injects object dynamics into the UNet by encoding sparse trajectories into multi-scale motion features; the Spatial Triple-Attention Transformer integrates image, text, and lighting information via parallel cross-attention modules. The model enables I2V generation conditioned on arbitrary combination of camera motion, object motion, and lighting direction.
We present VidCRAFT3, the first framework for high-fidelity Image-to-Video (I2V) generation with disentangled control over camera motion, object motion, and lighting direction. Existing methods struggle to simultaneously manage these elements due to entangled representations and insufficient annotated training data. To address this challenge, as shown in Fig. 2, our framework integrates three core components: (1) Image2Cloud, which reconstructs a 3D point cloud from a single input image and generates point cloud renderings along user-defined camera motion trajectory for precise camera control; (2) ObjMotionNet, which encodes sparse object trajectories into multi-scale motion features for realistic object dynamics; and (3) the Spatial Triple-Attention Transformer, which integrates lighting direction, image, and text features via parallel cross-attention layers. To overcome the scarcity of real-world videos with multi-element annotations, we construct three specialized datasets and adopt a three-stage training strategy to progressively optimize the model.
Our model is built upon the open-source I2V model DynamiCrafter, which consists of a VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ for image compression, a UNet with spatial and temporal layers, and CLIP text and image encoders. We employ a dual-stream injection approach where the reference image is encoded by CLIP image encoder and injected into UNet through image cross-attention. Simultaneously, the first frame of the point cloud rendering is replaced with the reference image, which is then encoded by the VAE encoder and concatenated with noise before being fed into the UNet. This design ensures seamless integration of spatial and temporal information for high-quality video generation.
3.2 Model Architecture
Camera Motion Control via Point Cloud Rendering. To achieve precise camera motion control in I2V, VidCRAFT3 leverages the Image2Cloud module, which reconstructs a high-quality 3D point cloud of the scene from a single reference image. Specifically, we employ DUSt3R, an unconstrained stereo 3D reconstruction model, to generate a 3D point cloud. Given a reference image $I_{\text{ref}}$, DUSt3R performs monocular or binocular reconstruction via point regression, followed by global alignment to ensure multi-view consistency: $\mathcal{P}=\text{DUSt3R}(I_{\text{ref}})$. The reconstructed point cloud provides explicit 3D geometry, enabling accurate rendering of the scene from arbitrary camera motion trajectory $E=\{E_{1},E_{2},\ldots,E_{F}\}$, where $F$ is the length of trajectory. Given a user-defined trajectory $E$, the point cloud rendering at frame $t$ is computed as $R_{t}=\pi(\mathcal{P},E_{t})$, where $\pi(\cdot)$ is the differentiable rendering function. However, due to the limitations of point cloud representation and the sparse 3D cues from a single image, the point cloud renderings may exhibit artifacts such as missing regions, occlusions, and geometric distortions. To address this, VidCRAFT3 integrates point cloud renderings as a input of VDM, which refines the coarse renderings to generate high-quality and temporally consistent video frames. This combination of explicit 3D geometry and generative refinement ensures both accurate camera control and realistic video synthesis.
Object Motion Control through Trajectory Learning. Object motion in VidCRAFT3 is controlled through sparse spatial trajectories. For up to $N$ objects in a $F$-frames video, each trajectory is defined as a sequence of 2D pixel coordinates $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\mid n\in[1,N],\,f\in[1,F]\}$, where $\mathbf{s}_{n}^{f}$ denotes the position of the $n$-th object in frame $f$. To model motion dynamics, we compute inter-frame optical flow vectors between consecutive frames. For each trajectory point $\mathbf{s}_{n}^{f}$, the displacement vector $\mathbf{v}_{n}^{f}$ is calculated as $\mathbf{v}_{n}^{f}=\mathbf{s}_{n}^{f+1}-\mathbf{s}_{n}^{f}=(x_{n}^{f+1}-x_{n}^{f},\,y_{n}^{f+1}-y_{n}^{f}).$ These sparse motion vectors are then projected into a per-frame optical flow map $\mathcal{V}^{f}\in\mathbb{R}^{H\times W\times 2}$, where $H$ and $W$ denote spatial dimensions. The mapping is formalized as
$\mathcal{V}^{f}(x,y)=\begin{cases}\mathbf{v}_{n}^{f}&\text{if }(x,y)=(x_{n}^{f},y_{n}^{f})\text{ for any }n,\\ (0,0)&\text{otherwise},\end{cases}$ (1)
with the first frame’s flow initialized as $\mathcal{V}^{1}(x,y)=(0,0),\forall(x,y)$. The full spatiotemporal flow tensor $\mathcal{V}\in\mathbb{R}^{F\times H\times W\times 2}$ is subsequently processed through Gaussian smoothing to obtain a dense motion representation $\tilde{\mathcal{V}}$. The ObjMotionNet is a neural network composed of multiple convolutional layers and downsampling operations, designed to extract multi-scale motion features from $\tilde{\mathcal{V}}$. Inspired by T2I-Adapter [33], ObjMotionNet injects multi-scale motion features exclusively into the UNet Encoder. This balances precise motion control with video quality, as the encoder handles structure while the decoder refines details, ensuring accurate guidance without compromising output quality.
Figure 3: Construction pipeline of the Object Motion Control Dataset.
Lighting Direction Control with Spatial Triple-Attention Transformer. The lighting direction is represented as a per-frame 3D vector $L=(l_{x},l_{y},l_{z})$, describing the orientation of light source in Cartesian coordinates. To effectively encode this directional information into a high-dimensional feature space, we employ Spherical Harmonic (SH) Encoding. SH encoding captures the angular characteristics of the lighting using basis functions up to degree 4, resulting in 16 coefficients. The resulting SH-encoded vector $L_{\text{SH}}\in\mathbb{R}^{16}$ is projected into the feature space of the UNet using a multi-layer perceptron (MLP). $\mathbf{E}_{\text{light}}=\text{MLP}(L_{\text{SH}})$, where $\mathbf{E}_{\text{light}}$ is a lighting embedding aligned with the dimensionality of text embedding.
To incorporate the lighting embedding into the UNet, we propose the Spatial Triple-Attention Transformer, which integrates three parallel attention modules: image cross-attention, text cross-attention, and lighting cross-attention. The lighting cross-attention module integrates the encoded lighting embedding $\mathbf{E}_{\text{light}}$ into the UNet. This attention mechanism modulates the spatial features based on the input lighting direction. The operation is defined as
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
where $Q$ (Query) comes from the self-attention output of the UNet, $K$ (Key) and $V$ (Value) are derived from $\mathbf{E}_{\text{light}}$. The outputs of three cross-attention are summed to produce a fused feature representation $\mathbf{O}=\mathbf{O}_{\text{image}}+\mathbf{O}_{\text{text}}+\mathbf{O}_{\text{light}}$ where $\mathbf{O}_{\text{image}}$, $\mathbf{O}_{\text{text}}$, and $\mathbf{O}_{\text{light}}$ are the outputs of the image, text, and lighting cross-attention modules, respectively. This novel mechanism ensures the generated videos maintain consistency across lighting, text, and image conditions.
Figure 4: Construction pipeline of the VLD Dataset
3.3 Dataset Construction
Due to the lack of datasets with annotations for camera motion trajectory, object motion trajectories, and lighting directions, we construct three specialized datasets. All datasets consist of 25-frame video clips with a spatial resolution of $320\times 512$ pixels.
Camera Motion Control Dataset. We constructe this dataset from RealEstate10K [75], creating 62,000 clips with smooth camera transitions. Using DUSt3R, we generated globally aligned point clouds and rendered the first frame’s point cloud along the ground-truth camera trajectory. Since RealEstate10K lacks captions, we uniformly sampled 4 frames per clip and generated descriptions using Qwen2-VL-7B-Instruct [53]. Despite its focus on indoor scenes, the dataset provides diverse camera trajectories for fine-grained motion control.
Object Motion Control Dataset. We construct this dataset from WebVid-10M [2], consisting of 60,000 video clips. The dataset construction pipeline, illustrated in Fig. 3, consists of the following five steps: (1) Clip Filtering: Clips with abrupt scene changes are removed using PySceneDetect, and sequences with temporal intervals (1–16 frames) are sampled. To retain only clips with significant motion, optical flow is computed via MemFlow [10], and the bottom 25% of clips with low motion scores are filtered out. (2) Dense Trajectory Generation: For each filtered clip, CoTrackerV3 [24] generates dense trajectories on a $16\times 16$ grid, capturing pixel-level motion dynamics across frames. (3) Trajectory Filtering by Length: Trajectories shorter than the average length are discarded to ensure meaningful object motion. (4) Sparse Trajectory Sampling: From the filtered dense trajectories, 1–8 sparse trajectories are sampled with probability proportional to their length. (5) Optical Flow Smoothing: Optical flow between adjacent frames is computed to encode motion direction and intensity. Finally, a Gaussian filter is applied to smooth the sparse trajectory matrix, ensuring stable training.
Lighting Direction Control Dataset. Since it is challenging to control a camera to follow the same trajectory under varying lighting conditions in real-world scenarios, collecting such datasets is both extremely difficult and costly. To address this, we introduce the VideoLightingDirection (VLD) Dataset, a novel synthetic dataset designed for modeling complex light-object interactions. The VLD Dataset comprises 57,600 synthetic videos rendered using Blender, featuring 3,600 realistic scenes, each with 16 videos captured under distinct lighting conditions. As illustrated in Fig. 4, the dataset construction pipeline consists of the following four steps: (1) Scene Creation: We designed two scene types: Haven scenes and BOP scenes. Haven scenes places 3D models from Poly Haven at the center of HDR environments, while BOP scenes randomly places BOP models in a six-plane textured room to simulate indirect lighting. Models and environments can be randomly combined for diversity. See supplementary materials for BOP scenes examples. (2) Camera Trajectory Sampling: We sample the starting point of the camera trajectory on a spherical region (radius: 0.7–1.3 meters) centered around the 3D models. A smooth trajectory is then randomly generated around the model, ensuring that the camera always facing the center of the model. (3) Lighting Direction Sampling: To enhance the lighting effect, we reduce the HDR environment intensity by 40%. We uniformly sample 16 points on a hemisphere centered on the model, with the base surface normal aligned with the camera’s viewing direction. Each sampled point serves as the position of a 2kW spotlight (radius: 1), oriented toward the model’s center. The light source position vectors are mapped to all camera coordinate systems and normalized to compute the lighting direction for each view. (4) Rendering and Annotation: Using Blender Cycles, we render each scene under 16 different lighting conditions while maintaining a consistent camera trajectory. Each frame is annotated with the corresponding camera pose and lighting direction, providing precise control signals for training.
3.4 Training Strategy
To enable simultaneous control over camera motion, object motion, and lighting direction, we adopt a three-stage progressive training strategy, designed to address the absence of datasets annotated with all three control signals.
Stage 1: Camera Motion Control Training. The training begins by initializing the model with DynamiCrafter pre-trained weights. The model is then fine-tuned on camera motion control dataset for 40,000 iterations, optimizing the entire UNet to align VDM with camera motion. This stage establishes robust 3D scene understanding by integrating point cloud renderings, ensuring precise global camera movements while maintaining temporal consistency.
Stage 2: Dense Object Trajectories and Lighting Mixed Fine-tuning. We combine the Object Motion Control Dataset and VLD Datasets to create a comprehensive dataset annotated with camera motion, object motion, and lighting direction. This hybrid dataset enhances the ability of model to learn joint control over these three conditions. Dense object trajectories are incorporated to provide rich motion details, accelerating model convergence. To retain the camera control capabilities from stage 1, the temporal layers of the UNet remain frozen, while the spatial layers and newly introduced components, including ObjMotionNet, lighting cross-attention, and the MLP for lighting direction projection, are optimized for 20,000 iterations. This stage enables the model to simultaneously control all three conditions while ensuring global camera alignment.
Stage 3: Sparse Object Trajectories and Lighting Mixed Fine-tuning. The same hybrid dataset from Stage 2 is reused, but dense trajectories are replaced with sparse trajectories to simulate real-world user interactions. Fine-tuning follows the same parameters as Stage 2, with 20,000 iterations. This stage forces the model to infer complex motion patterns from limited trajectory data, and the progressive shift from dense to sparse supervision enhances its ability to generalize to practical scenarios. By progressively learning camera motion, object motion, and lighting direction controls across stages while strategically freezing layers to retain prior knowledge, the training strategy ensures that VidCRAFT3 achieves fine-grained, synergistic control over all three elements, even in the absence of fully annotated multi-task training data.
4 Experiments
Figure 5: Qualitative comparisons with SOTA methods on RealEstate10K.
Figure 6: Qualitative comparisons with SOTA methods on WebVid-10M.
4.1 Experimental Setup
Implementation Details. Our model builds upon DynamiCrafter, initialized with its pre-trained weights. During training and inference, video clips are processed at $320\times 512$ resolution with 25 frames. To optimize the model, we employ the Adam optimizer with a learning rate of $1\times 10^{-5}$ and a batch size of 96. The training is conducted on 8 NVIDIA H100 GPUs. For inference, we adopt the DDIM sampler and classifier-free guidance to enhance generation quality and multimodal consistency. During inference, the model averagely uses 20 GB of GPU memory and takes 42 seconds.
[…]
CameraCtrl 97.99 96.11 29.41 4.19
MotionCtrl 103.82 188.93 30.18 4.23
CamI2V 98.54 85.03 30.37 4.24
Ours 75.62 49.77 32.32 4.07
Table 1: Quantitative comparison with SOTA methods on RealEstate10K.
Evaluation Datasets. We evaluate our model on three domain-specific datasets and a generalized test set. For camera motion control, we sample 1,000 samples from the RealEstate10K test set. For object motion control, we sample 1,000 samples from the WebVid-10M test set. For lighting direction control, we sample 1,000 samples from the VLD dataset, covering a wide range of lighting directions. To provide a broader evaluation, we create a generalized test set consisting of 100 videos sourced from copyright-free websites like Pixabay and Pexels, as well as videos generated by T2V models. This dataset spans categories such as human activities, animals, vehicles, indoor scenes, artworks, natural landscapes, and AI-generated images.
Evaluation Metrics. We evaluate VidCRAFT3 across three dimensions: (1) Video Quality: FID for assessing visual fidelity, FVD for measuring temporal coherence, and CLIPSIM for detecting semantic alignment; (2) Motion Control Performance: Based on camera poses estimated by DUSt3R and object trajectories extracted by CoTrackerV3, motion control performance is quantified using CamMC and ObjMC metrics [57], which measure the Euclidean distance between predicted and ground-truth values; (3) Lighting Control Effectiveness: Evaluated through LPIPS, SSIM, and PSNR, comparing generated frames with ground-truth images to assess perceptual quality and structural fidelity.
4.2 Comparisons with State-of-the-Art Methods
Due to the lack of open-source I2V methods capable of simultaneously controlling camera and object motion, we evaluate VidCRAFT3 separately for camera motion control and object motion control, comparing it against SOTA methods in each domain.
[…]
Image Conductor 150.26 242.01 29.69 12.96
Motion-I2V 128.35 171.35 30.92 3.96
Ours 87.12 120.65 32.99 3.51
Table 2: Quantitative comparison with SOTA methods on WebVid-10M.
Camera Motion Control. VidCRAFT3 demonstrates superior performance in camera motion control compared to state-of-the-art (SOTA) methods like CameraCtrl, CamI2V, and MotionCtrl on the RealEstate 10K dataset. As shown in Table 1 and Fig. 5, VidCRAFT3 excels in camera motion control, as evidenced by its CamMC score of 4.07, which is lower than CameraCtrl (4.19), CamI2V (4.24), and MotionCtrl (4.23). VidCRAFT3 also achieves better results in other key metrics indicate that VidCRAFT3 not only controls camera motion more accurately but also generates videos with higher visual quality, temporal coherence, and semantic alignment. Qualitatively, VidCRAFT3 generates smoother and more realistic camera movements with fewer artifacts, particularly in complex scenes. This success stems from the Image2Cloud module, which enable precise 3D scene reconstruction and fine-grained camera control.
Object Motion Control. VidCRAFT3 demonstrates exceptional performance in object motion control, as highlighted in Table 2. The model achieves an ObjMC score of 3.51, which is lower than Image Conductor (12.96) and Motion-I2V (3.96). This indicates that VidCRAFT3 more accurately aligns the generated object trajectories with the ground truth, resulting in more realistic and faithful object motion. Additionally, VidCRAFT3 outperforms in other key metrics, showcasing superior visual quality, temporal coherence, and semantic alignment. Qualitatively, as shown in Fig. 6, VidCRAFT3 generates more realistic and consistent object movements compared to Image Conductor and Motion-I2V. The model effectively captures the dynamics of object motion, ensuring smooth transitions and natural interactions within the scene. These results highlight VidCRAFT3’s robust object motion control, driven by its advanced ObjMotionNet, which effectively captures and controls complex motion patterns.
[…]
CameraCtrl 3.4% 5.4% 4.7%
MotionCtrl 5.7% 7.3% 6.5%
CamI2V 5.6% 6.8% 5.2%
Ours 85.3% 80.5% 83.6%
[…]
Image Conductor 10.3% 11.9% 11.2%
Motion-I2V 14.8% 10.9% 13.2%
Ours 74.9% 77.2% 75.6%
Table 3: User study. Results demonstrating our method’s superior performance in camera and object motion control compared to baseline approaches across all metrics.
[…]
Text Cross-Atten 111.08 121.95 22.77 18.14 0.72 0.13 5.31
Time Embed. 101.71 123.31 22.70 19.07 0.73 0.12 5.21
Lighting Cross-Attn 100.83 117.69 23.70 19.49 0.74 0.11 5.00
Table 4: Ablation of lighting embedding integration strategies on VLD. Text Cross-Atten method concatenates lighting embedding with text embedding and integrates them into the model through text cross-attention. Time Embed method adds lighting embedding to time embedding. The proposed Lighting Cross-Atten method, detailed in Sec. 3.2, demonstrates enhanced performance in integrating lighting conditions.
User Study. To evaluate our method, we conducted a user study with 10 participants assessing 200 randomly selected results from our method and five baselines. The participants evaluated the results based on four key metrics: Camera Motion Precision, Object Motion Precision, Visual Quality, and Overall Quality. As shown in Table 3, our method achieved over 80% in camera motion control and over 74% in object motion control across all metrics, demonstrating precise and visually appealing control, validating its effectiveness and robustness in real-world applications.
4.3 Ablation Study
Training Strategy. We compare training with Dense, Sparse, and Dense+Sparse trajectories. As shown in Table 5, Dense training (ObjMC: 4.39) provides rich motion information but struggles during inference due to sparse inputs. Sparse training (ObjMC: 4.05) improves inference alignment but lacks motion detail. The Dense+Sparse approach achieves the best results (ObjMC: 3.51), leveraging dense trajectories for robust learning and fine-tuning with sparse trajectories for better generalization. This hybrid strategy enhances object motion control and video quality, as shown by superior FID (87.12) and FVD (120.65) scores.
Lighting Embedding Integration Strategies. We compare different lighting embedding integration strategies. As shown in Table 4, Lighting Cross-Atten achieves the best results, with an PSNR of 19.49, SSIM of 0.74, and LPIPS of 0.11, outperforming other methods. These findings confirm the effectiveness of Lighting Cross-Attention for precise lighting direction control.
Representation of Lighting Direction. We compare Fourier Embedding [31] and SH Encoding for representation of lighting direction. As shown in Table 6, SH Encoding outperforms Fourier Embedding, achieving better PSNR (19.49), SSIM (0.74), and LPIPS (0.11). These results confirm SH Encoding’s effectiveness for precise lighting control in image-to-video generation. For more results, please refer to the supplementary materials.
[…]
Dense 92.05 143.44 30.78 4.39
Sparse 91.54 123.15 30.93 4.05
Dense+Sparse 87.12 120.65 32.99 3.51
Table 5: Ablation of training strategy on WebVid-10M.
[…]
Fourier Embedding 107.40 121.89 21.71 17.48 0.70 0.14 5.03
SH Encoding 100.83 117.69 23.70 19.49 0.74 0.11 5.00
Table 6: Ablation of representation of lighting direction on VLD.
5 Conclusions
We present VidCRAFT3, a novel framework for precise image-to-video generation with simultaneous control over camera motion, object motion, and lighting direction. It integrates Image2Cloud for 3D point cloud generation, ObjMotionNet for encoding sparse object trajectories, and a Spatial Triple-Attention Transformer for incorporating lighting direction. The VideoLightingDirection dataset and a three-stage training strategy address real-world annotation challenges. Experiments show that VidCRAFT3 outperforms state-of-the-art methods in control granularity and visual coherence, advancing video generation toward more realistic and versatile applications.
Supplementary Material
Appendix A Preliminary
Image-to-video generation. Given a reference image $I_{\text{ref}}$ and some text description, image-to-video generation aims to generate a sequence of video frames $I=\{I_{1},I_{2},\ldots,I_{F}\}$ with $I_{1}=I$. In this work, we also allow the user to provide three forms of control signals: camera motion, object motion, and lighting direction. The video frames will be generated according to the text description and all the control signals provided. Specifically, the camera motion is represented by the movement of a set of camera extrinsics $E=\{E_{1},E_{2},\ldots,E_{F}\}$; the object motion is represented by a few sparse sequences of image pixels $\{s_{1}^{1},s_{1}^{2},s_{1}^{3},\ldots,s_{1}^{F}\},\{s_{1}^{2},s_{2}^{2},s_{2}^{3},\ldots,s_{2}^{F}\},\ldots$; the lighting direction is simply a unit 3d vector $L\in\boldsymbol{R}^{3}$.
Video Diffusion model (VDMs) represent a class of generative models that extend the principles of image diffusion to the domain of video generation. These models operate by defining a forward diffusion process that gradually transforms an initial video sample $x_{0}\sim p_{\text{data}}(x)$ into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ timesteps. The reverse process, parameterized by a denoising network $\epsilon_{\theta}(x_{t},t)$, learns to iteratively denoise the noisy latent representation $x_{t}$ to recover the original data $x_{0}$. The training objective is formulated as:
$\min_{\theta}\mathbb{E}_{t,x\sim p_{\text{data}},\epsilon\sim\mathcal{N}(0,I)}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2},$
where $\epsilon$ represents the ground truth noise, and $\theta$ denotes the learnable parameters of the network. Once trained, the model can generate high-quality videos by sampling from a random noise distribution $x_{T}$ and applying the learned denoising process iteratively.
To address the computational challenges associated with high-dimensional video data, Latent Diffusion Models (LDMs) are often employed. In this framework, a video $x\in\mathbb{R}^{F\times 3\times H\times W}$ is first encoded into a lower-dimensional latent space $\mathbf{z}=\mathcal{E}(x)$, where $\mathbf{z}\in\mathbb{R}^{F\times C\times h\times w}$. The diffusion and denoising processes are then performed in this latent space, significantly reducing computational complexity. The denoising process is conditioned on additional inputs $\mathbf{c}$, such as text prompts or motion control signals, enabling the generation of videos that adhere to specific semantic or temporal constraints. The final video is reconstructed through a decoder $\hat{x}=\mathcal{D}(\mathbf{z})$, ensuring both spatial and temporal coherence in the generated output.
Appendix B VLD Dataset Samples
Fig. 7 illustrates representative samples from our proposed VideoLightingDirection (VLD) Dataset, a synthetic dataset specifically designed to model complex light-object interactions. The dataset comprises two types of scenes: (a) Haven scenes, in which 3D models from Poly Haven are placed centrally within HDR environments; and (b) BOP scenes, featuring randomly positioned BOP models in six-plane textured rooms to simulate indirect lighting. Each subset includes video frames captured under two distinct lighting conditions while maintaining a consistent camera trajectory. This design highlights how changes in lighting direction impact object shading, reflections, and overall visual coherence, providing valuable data for training and evaluating models requiring precise lighting direction annotations.
Figure 7: Illustrations of samples from the proposed VideoLightingDirection (VLD) Dataset, featuring synthetic scenes designed to model complex light-object interactions. (a) Haven scenes, where 3D models from Poly Haven are placed at the center of HDR environments. (b) BOP scenes, with BOP models randomly positioned within six-plane textured rooms to simulate indirect lighting. Each subset includes video frames captured under two distinct lighting conditions, maintaining consistent camera trajectories.
Appendix C Qualitative Results of Ablation Study
In this section, we present qualitative results of the ablation studies described in the main text, providing visual insights into how different design choices impact the generated videos.
Training Strategy. Fig. 8 qualitatively compares videos generated using Dense, Sparse, and Dense+Sparse training strategies. Dense training results exhibit detailed object motion but struggle with precise alignment when tested with sparse trajectories. Sparse training achieves better alignment but at the cost of reduced motion detail. The Dense+Sparse approach effectively combines the strengths of both, delivering superior motion coherence and overall visual quality.
Lighting Embedding Integration Strategies. Fig. 9 compares the generated videos of different lighting embedding integration strategies. Lighting Cross-Attention demonstrates clear superiority, providing precise control over lighting directions, resulting in more realistic shadows, reflections, and overall lighting effects compared to Text Cross-Attention and Time Embedding.
Representation of Lighting Direction. Fig. 10 presents a visual comparison between Fourier Embedding and SH Encoding for representing lighting direction. SH Encoding consistently produces more realistic and detailed lighting effects, including enhanced shading and reflections, further supporting the quantitative findings in the main text.
These qualitative analyses further validate our key design decisions, emphasizing their impact on generating precise and realistic image-to-video translations under controlled camera motion, object motion, and lighting direction.
Appendix D Additional Qualitative Results
This section provides additional qualitative comparisons illustrating VidCRAFT3’s capabilities across multiple control dimensions.
Camera Motion + Lighting Direction Control. Fig. 11 demonstrates VidCRAFT3’s effectiveness in simultaneously controlling camera motion and lighting direction, showing clear advantages in lighting realism and trajectory accuracy compared to original sequences.
Camera Motion + Object Motion Control. Fig. 12 presents examples of simultaneous control over camera and object motions. VidCRAFT3 accurately reproduces intended object trajectories and maintains specified camera movements, closely matching the ground truth.
Camera Motion Control. Fig. 13 compares VidCRAFT3’s results with state-of-the-art methods (CameraCtrl, MotionCtrl, CamI2V) and ground-truth sequences, highlighting superior camera trajectory accuracy and visual coherence.
Object Motion Control. Fig. 14 provides comparisons against Image Conductor and Motion-I2V, showcasing VidCRAFT3’s superior capability in generating realistic object trajectories and visually coherent motion across diverse scenarios.
Lighting Direction Control. Fig. 15 further illustrates VidCRAFT3’s precise control over lighting direction, achieving accurate shading, shadows, and reflections, consistent with ground-truth sequences.
These additional qualitative evaluations highlight VidCRAFT3’s comprehensive and precise control capabilities, confirming its robustness and versatility across different control dimensions.
Appendix E Limitations
VidCRAFT3 can generate unstable results for many specific situations, e.g., large human motion, phyisical interactions, significant change of lighting conditions. This is mostly due to the lack of diverse training data and we believe the neural architecture can also be improved to promote a better understanding of physics and 3D spatial relationships of objects. The camera motion control and object motion control datasets contain inaccurate annotations of camera pose and motion trajectories, which sometimes result in blurred generated videos. Currently, VidCRAFT3 only offers control over the lighting direction and therefore can be improved to support full HDR or other more fine-grained representations of the light field.
Figure 8: Qualitative comparisons of training strategies (Dense, Sparse, Dense+Sparse) in the ablation study conducted on WebVid-10M. Results clearly illustrate that Dense training struggles with alignment given sparse input trajectories during inference, Sparse training improves alignment but loses motion detail, while Dense+Sparse training effectively combines both strategies, resulting in superior alignment and enhanced motion realism compared to the ground truth (GT).
Figure 9: Qualitative comparison of lighting embedding integration strategies (Text Cross-Attn, Time Embedding, Lighting Cross-Attn) in the ablation study on the VLD dataset. Results clearly demonstrate that the Lighting Cross-Attn strategy provides more accurate and realistic lighting control, achieving superior shadows, reflections, and overall visual fidelity compared to Text Cross-Attn and Time Embedding, aligning closely with the ground truth (GT).
Figure 10: Qualitative comparison between Fourier Embedding and SH Encoding for representing lighting direction on the VLD dataset. Results demonstrate that SH Encoding consistently produces more realistic lighting effects, achieving better shading accuracy and reflection details compared to Fourier Embedding, aligning closely with the ground truth (GT).
Figure 11: Additional qualitative results demonstrating simultaneous control over camera motion and lighting direction. Each pair of rows compares the original video with the corresponding relighting results generated by VidCRAFT3, highlighting the framework’s effectiveness in precisely adjusting lighting direction while preserving accurate camera motion.
Figure 12: Additional qualitative results showcasing simultaneous camera motion and object motion control. Each example compares our generated video (Ours) with ground-truth videos (GT), demonstrating VidCRAFT3’s capability to accurately reproduce intended object motion while precisely maintaining the specified camera motion.
Figure 13: Additional qualitative comparisons for camera motion control. Results from VidCRAFT3 (Ours) are compared with state-of-the-art methods (CameraCtrl, MotionCtrl, CamI2V) and ground-truth (GT) videos. These examples demonstrate VidCRAFT3’s superior capability in generating visually coherent and precise camera motion across diverse scenarios.
Figure 14: Additional qualitative comparisons for object motion control. Results generated by VidCRAFT3 (Ours) are compared against state-of-the-art methods (Image Conductor, Motion-I2V) and ground-truth (GT) videos. The examples demonstrate VidCRAFT3’s improved capability in accurately reproducing specified object trajectories, achieving more realistic and coherent object motion across diverse scenarios.
Figure 15: Additional qualitative results demonstrating lighting direction control. Our generated videos (”Ours”) are compared against ground-truth sequences (”GT”), clearly highlighting VidCRAFT3’s capability to accurately reproduce different lighting direction, achieving realistic shading, shadowing, and reflection details across diverse scenes.

### v4

@@
 Abstract
-Recent image-to-video generation methods have demonstrated success in enabling control over one or two visual elements, such as camera motion or object motion.
-However, these methods are unable to offer control over multiple visual elements due to limitations in data and network efficacy.
-In this paper, we introduce VidCRAFT3, a novel framework for precise image-to-video generation that enables control over camera motion, object motion, and lighting direction simultaneously.
-VidCRAFT3 integrates three core components: Image2Cloud generates 3D point cloud from a reference image; ObjMotionNet encodes sparse object trajectories using multi-scale optical flow features; and Spatial Triple-Attention Transformer incorporates lighting direction embeddings via parallel cross-attention modules.
-Additionally, we introduce the VideoLightingDirection dataset, providing synthetic yet realistic video clips with accurate per-frame lighting direction annotations, effectively mitigating the lack of annotated real-world datasets.
-We further adopt a three-stage training strategy, ensuring robust learning even without joint multi-element annotations.
-Extensive experiments show that VidCRAFT3 produces high-quality video content, outperforming state-of-the-art methods in control granularity and visual coherence.
-Code and data will be publicly available.
+Controllable image-to-video (I2V) generation transforms a reference image into a coherent video guided by user-specified control signals.
+In content creation workflows, precise and simultaneous control over camera motion, object motion, and lighting direction enhances both accuracy and flexibility.
+However, existing approaches typically treat these control signals separately, largely due to the scarcity of datasets with high-quality joint annotations and mismatched control spaces across modalities.
+We present VidCRAFT3, a unified and flexible I2V framework that supports both independent and joint control over camera motion, object motion, and lighting direction by integrating three core components.
+Image2Cloud reconstructs a 3D point cloud from the reference image to enable precise camera motion control.
+ObjMotionNet encodes sparse object trajectories into multi-scale optical flow features to guide object motion.
+The Spatial Triple-Attention Transformer integrates lighting direction embeddings via parallel cross-attention.
+To address the scarcity of jointly annotated data, we curate the VideoLightingDirection (VLD) dataset of synthetic static-scene video clips with per-frame lighting-direction labels, and adopt a three-stage training strategy that enables robust learning without fully joint annotations.
+Extensive experiments show that VidCRAFT3 outperforms existing methods in control precision and visual coherence.
+Code and data will be released.
+Project page: https://sixiaozheng.github.io/VidCRAFT3/.
+publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
+I Introduction
 Figure 1: VidCRAFT3 is the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction.
@@
 VidCRAFT3 can take any combination of supported control signals and deliver fine-grained and faithful generation results.
-1 Introduction
-Image-to-video (I2V) generation is a powerful technique that brings still images to life, enabling widespread applications in the fields of content creation, advertising, and art animation.
-Typically, I2V generation aims to animate a reference image according to user-defined control signals, such as text, object motion, and camera motion, while maintaining the quality and natural dynamics of the generated video.
-Driven by advances in diffusion-based generative models and extensive web-scale data [2], recent methods have significantly enhanced the ability to generate temporally coherent and visually compelling videos from limited inputs such as a single image or sparse annotations [15, 6, 60, 14, 57, 17, 28, 13].
-However, achieving precise and simultaneous control over multiple video attributes, such as camera motion, object motion, and lighting direction, remains a critical yet challenging objective.
-Existing approaches typically address these control attributes independently or in a partially integrated manner.
-Methods focusing exclusively on camera control, such as CameraCtrl [17] and ViewCrafter, do not adequately handle detailed object motion or dynamic lighting variations.
-Similarly, object-motion-centric frameworks, including DragAnything [58], lack integrated capabilities for precise camera and lighting control.
-For lighting control, techniques like DILightNet [66] and NeuLighting [27] provide lighting control but often lack generalizability due to reliance on specific categories (e.g., human faces) and object-specific annotations.
-Other methods [27, 68, 22, 68, 22] also primarily target narrow domains, such as human faces, further restricting their applicability to open-domain scenarios.
-Simultaneous and fine-grained control of camera motion, object motion, and lighting direction introduces multiple significant challenges: (1) Accurate camera motion control from single-image inputs requires robust extraction of sufficient 3D geometric information.
-(2) Realistic and detailed object motion control demands effective representation and refinement of sparse motion trajectories without compromising visual fidelity.
+Image-to-video (I2V) generation is a powerful technique that brings still images to life, with broad applications across content creation, advertising, and animation.
+Controllable I2V generation, in particular, aims to animate a reference image according to user-specified control signals, such as text, object motion, and camera motion, while preserving high visual fidelity.
+Recent advancements in diffusion-based generative models and extensive web-scale data [1] have significantly enhanced the ability to generate temporally coherent and visually compelling videos from limited inputs, such as a single image or sparse annotations [2, 3, 4, 5, 6, 7].
+However, achieving precise, simultaneous control of camera motion, object motion, and lighting direction remains a significant challenge in real-world scenarios.
+Existing approaches typically address these control signals independently or in a partially integrated manner.
+Methods focusing exclusively on camera control, such as CameraCtrl [7] and CamTrol [8], fail to adequately handle detailed object motion or dynamic lighting variations.
+Similarly, object-motion-centric frameworks like DragAnything [9], lack integrated capabilities for precise camera and lighting control.
+For lighting control, techniques like DILightNet [10] and NeuLighting [11] enable lighting adjustments.
+However, these methods often lack generalizability due to their reliance on specific categories, such as human faces, and the use of HDR maps.
+In addition, the dependence on control signals such as HDR maps [12] and background videos [13] restricts interactive control of lighting direction, making these methods less in open-domain scenarios.
+Simultaneous and precise control of camera motion, object motion, and lighting direction introduces multiple significant challenges: (1) Accurate camera motion control from a reference image requires reliable 3D priors to prevent drift and geometric distortion under large viewpoint changes.
+(2) Realistic and detailed object motion control demands effective representation of sparse object trajectories (e.g., a sequence of 2D keypoints) without compromising visual fidelity.
 (3) Dynamic lighting control necessitates integrating illumination adjustments coherently with both camera and object motion to ensure temporal consistency and visual realism.
+(4) High-quality datasets with joint annotations for camera motion, object motion, and lighting direction remain scarce.
 To overcome these challenges, we propose VidCRAFT3, which integrates three core components.
-First, the Image2Cloud module leverages DUSt3R [54] to reconstruct a 3D point cloud from a single reference image, enabling precise camera motion control by rendering the point cloud along user-defined camera motion trajectory.
-Second, ObjMotionNet encodes sparse object motion trajectories by extracting multi-scale motion features from Gaussian-smoothed optical flow maps to guide realistic object motion.
-Third, Spatial Triple-Attention Transformer integrates lighting direction control by integrating lighting embedding with image-text embeddings through parallel cross-attention layers, ensuring consistent illumination effects.
-The capabilities of these modules are shown in Fig.
-1, showcasing VidCRAFT3’s control over camera motion, object motion, and lighting direction.
-To address data scarcity, we introduce the VideoLightingDirection (VLD) dataset, providing synthetic yet highly realistic video sequences with accurate per-frame lighting annotations.
-Additionally, we develop a three-stage training strategy that enables effective learning without joint multi-element annotations.
-The main contributions of this paper are: (1) VidCRAFT3 is the first I2V framework to achieve simultaneous control over camera motion, object motion, and lighting direction through a disentangled architecture combining Image2Cloud, ObjMotionNet, and Spatial Triple-Attention Transformer.
-(2) VLD dataset provides synthetic videos accompanied by accurate per-frame lighting direction annotations, effectively addressing the scarcity of annotated real-world datasets.
-(3) A three-stage training strategy enables robust multi-element control without requiring real-world videos annotated with camera, object, and lighting signals.
+First, the Image2Cloud module leverages DUSt3R [14] to reconstruct a 3D point cloud from a single reference image, enabling precise camera motion control by rendering the point cloud along user-defined camera trajectory.
+Second, ObjMotionNet encodes sparse object trajectories by extracting multi-scale motion features from Gaussian-smoothed optical flow maps to guide realistic object motion.
+Third, Spatial Triple-Attention Transformer integrates lighting embedding with image-text embeddings through parallel cross-attention layers.
+Capabilities of these modules are illustrated in Fig.
+1, highlighting VidCRAFT3’s control over camera motion, object motion, and lighting direction.
+To address data scarcity, we introduce the VideoLightingDirection (VLD) dataset, providing synthetic static-scene video clips that are highly realistic, along with accurate per-frame lighting annotations.
+Additionally, we develop a three-stage training strategy that enables effective learning without fully joint annotations.
+The main contributions of this paper are: (1) To the best of our knowledge, VidCRAFT3 is the first I2V framework to achieve simultaneous control over camera motion, object motion, and lighting direction through a disentangled architecture combining Image2Cloud, ObjMotionNet, and Spatial Triple-Attention Transformer.
+(2) The VLD dataset provides synthetic videos with accurate per-frame lighting direction annotations, effectively addressing the scarcity of datasets with fully joint annotations.
+(3) A three-stage training strategy enables robust multi-element control without requiring dataset annotated with camera, object, and lighting signals.
 (4) Extensive experiments demonstrate that VidCRAFT3 achieves state-of-the-art performance surpassing existing methods in terms of control precision, visual quality, and generalization capabilities.
-2 Related Work
-2.1 Image-to-video Generation
+II Related Work
+II-A Image-to-video Generation
 Image-to-video (I2V) generation [2, 3, 15, 4, 5, 16] aims to animate static images into dynamic videos while preserving visual content and introducing realistic motion.
@@
 Another line of work, exemplified by SEINE [21], DynamiCrafter [4] and PixelDance [22], expands the input channels of diffusion models to concatenate the static image with noisy latents, effectively injecting image information into the model.
-However, these methods preserve input image fidelity while generating dynamic videos but often struggle with fine-grained details due to reliance on global conditions.
-2.2 Motion-controlled Video Generation
-Motion-controlled video generation [57, 17, 64] focuses on creating high-fidelity videos with user-defined motion dynamics.
-Existing methods can be categorized into camera motion control, object motion control, and joint motion control.
-For camera motion control [17, 1, 26, 21, 56], CamI2V [72], CameraCtrl [17], and CamCo [61] encode camera parameters into Plücker embeddings for more discriminative motion representation.
-ViewCrafter [65] and I2VControl-Camera [11] use 3D representations of reference images to render point clouds and generate partial frames.
-For object motion control [32, 46, 55, 23, 30, 71, 74, 62, 37, 19, 12, 34, 59], existing methods employ diverse strategies to guide object movements in video synthesis.
-DragNUWA [64], Image Conductor [28], DragAnything [58], and MotionBridge [50] utilize sparse optical flow to control object motion, while MOFA-Video [36] and Motion-I2V [45] predict dense motion fields to control object motion.
-Alternatively, Boximator [52] and Direct-a-Video [63] rely on bounding box trajectories to guide object motion.
-LeviTor [51] combines depth and K-means clustering for precise 3D trajectory control.
-Current research on Joint Motion Control [63, 57, 13, 9], aiming to simultaneously control camera and object motions, remains limited.
-MotionCtrl [57] designs specialized modules for precise control, while Motion Prompting [13] uses motion tracks for flexible guidance.
-Perception-as-Control [9] further enhances control with 3D-aware motion representations.
+However, these methods preserve input image fidelity during dynamic video generation but often struggle with fine-grained details due to reliance on global conditions.
+II-B Motion-controlled Video Generation
+Motion-controlled video generation [6, 23] focuses on creating high-fidelity videos with user-defined motion dynamics.
+Recent video generation models [24, 25, 26, 27, 28, 29] achieve impressive visual quality, but rely only on text and/or images for controlling content and motion, leading to coarse and implicit motion control.
+Existing approaches can be broadly divided into camera motion control, object motion control, and joint motion control.
+In the domain of camera motion control, one line of work conditions models on explicit camera intrinsics and extrinsics or on Plücker embeddings [7, 30, 31, 32, 33, 34, 35, 6, 36, 37].
+For instance, MotionCtrl [6] injects extrinsic matrices into temporal attention to modulate viewpoint, whereas CameraCtrl [7] utilizes Plücker embeddings to incorporate geometric structure.
+Although these methods show promising results, directly mapping camera parameters to generative dynamics often restricts precision and generalization, particularly for trajectories outside the training distribution.
+To improve geometric consistency, another line of research lifts a reference image into 3D using depth maps or point clouds [38, 8, 39, 40, 41, 42, 43, 44, 45, 46, 47].
+Representative examples include ViewCrafter [44] and CamTrol [8], which render partial frames from reconstructed point clouds as guidance during generation.
+A further research direction explores training-free solutions [8, 48, 49, 50, 51] that derive motion representations by inverting temporal attention maps in pretrained models, as exemplified by MotionMaster [48] and MotionClone [50].
+For object motion control, prior work explores complementary control signals and injection schemes.
+Flow-conditioned methods steer dynamics with sparse optical flow derived from user drags [52, 53, 54], as exemplified by DragNUWA [23], Image Conductor [55], and ReVideo [56].
+Dense motion fields further refine pixelwise trajectories in MOFA-Video [57] and Motion-I2V [58].
+Region-level cues provide a low-overhead interface, with bounding-box trajectories guiding objects[59, 60] in Boximator [61] and TrailBlazer [62], MagicMotion [63].
+Region-level cues offer a low-overhead interface, where bounding-box trajectories guide object motion [59, 60, 62, 64], as demonstrated by Boximator [61] and MagicMotion [63].
+To improve geometric fidelity, recent work leverages keypoint/point-map, mask-aware, and 3D cues.
+TrackGo [65] injects free-form masks and arrows via a lightweight temporal adapter.
+Tora [66] encodes trajectory maps with a 3D VAE.
+LeViTor [67] fuses depth with clustered keypoints for precise 3D control.
+Segmentation-aware schemes align user intent with object extent, with DragEntity [68] and DragAnything [9] mapping user drags to the corresponding masked entities.
+In parallel, training-free variants steer inference via attention guidance, energy-based denoising, or latent edits, enabling box/mask-conditioned control without finetuning [69, 70, 71].
+Current research on Joint Motion Control [72, 6, 73, 74, 75, 76, 77, 78], aiming to simultaneously control both camera and object motions, remains limited.
+Perception-as-Control [74] uses a simplified 3D scene to produce spatially aligned control signals; Motion Prompting [73] encodes spatiotemporal directives as structured prompts; MotionCanvas [77] converts 3D intent to 2D conditioning for diffusion models; CineMaster [78] combines 3D boxes and projected depth with a camera adapter; MotionAgent [76] decomposes text into camera extrinsics and object trajectories and composes them as optical flow.
 Unlike existing methods, we propose VidCRAFT3, the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction.
 By combining 3D point cloud rendering, trajectory learning, and Spatial Triple-Attention Transformer, our approach effectively decouples these elements, ensuring temporal consistency and enhanced realism in complex scenes.
-2.3 Lighting-controlled Visual Generation
+II-C Lighting-controlled Visual Generation
 Lighting-controlled visual generation aims to manipulate illumination while preserving scene geometry and materials.
 Previous methods primarily focus on portrait lighting [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89], laying the foundation for effective and accurate illumination modeling.
-Recent advances in diffusion models have significantly improved the quality and flexibility of lighting control.
+Recent advances in diffusion models significantly improve the quality and flexibility of lighting control.
 Methods like DiLightNet [10] and GenLit [90] achieve fine-grained and realistic relighting through radiance hints and SVD, respectively.
@@
 Recent techniques leverage 3D-aware generative models for temporally consistent relighting, as seen in EdgeRelight360 [92] and ReliTalk [93].
-Neural rendering approaches [5, 68] use datasets like dynamic OLAT for high-quality portrait video relighting, while reflectance field-based methods [22] infer lighting from exemplars.
-Despite these advancements, most prior works focus on portraits and HDR lighting conditions.
-In contrast, VidCRAFT3 targets general scenes, enabling interactive adjustment of lighting directions.
-3 Method
-3.1 Overview
+Neural rendering approaches [94, 95] use datasets like dynamic one-light-at-a-time (OLAT) for high-quality portrait video relighting, while reflectance field-based methods [96] infer lighting from exemplars.
+Light-A-Video [97] achieves training-free, temporally consistent video relighting via a Consistent Light Attention module and Progressive Light Fusion, while RelightVid [13] employs a 3D UNet with temporal layers trained on LightAtlas to deliver temporally consistent, high-quality results under flexible conditions (text, background videos, HDR maps) without intrinsic decomposition.
+Despite these advances, most prior work emphasizes portraits and HDR lighting conditions, limiting interactive control in general scenes.
+In contrast, VidCRAFT3 targets open, non-portrait scenarios and provides interactive, direction-level lighting control with geometry/material preservation and a scene-agnostic control interface.
+III Method
+III-A Overview
 Figure 2: Architecture of VidCRAFT3 for controlled image-to-video generation.
-The model builds on Video Diffusion Model (VDM) and consists of three main components: the Image2Cloud reconstructs 3D point cloud from a single reference image and generates point cloud renderings along user-defined camera motion trajectory; the ObjMotionNet injects object dynamics into the UNet by encoding sparse trajectories into multi-scale motion features; the Spatial Triple-Attention Transformer integrates image, text, and lighting information via parallel cross-attention modules.
-The model enables I2V generation conditioned on arbitrary combination of camera motion, object motion, and lighting direction.
-We present VidCRAFT3, the first framework for high-fidelity Image-to-Video (I2V) generation with disentangled control over camera motion, object motion, and lighting direction.
-Existing methods struggle to simultaneously manage these elements due to entangled representations and insufficient annotated training data.
-To address this challenge, as shown in Fig.
-2, our framework integrates three core components: (1) Image2Cloud, which reconstructs a 3D point cloud from a single input image and generates point cloud renderings along user-defined camera motion trajectory for precise camera control; (2) ObjMotionNet, which encodes sparse object trajectories into multi-scale motion features for realistic object dynamics; and (3) the Spatial Triple-Attention Transformer, which integrates lighting direction, image, and text features via parallel cross-attention layers.
-To overcome the scarcity of real-world videos with multi-element annotations, we construct three specialized datasets and adopt a three-stage training strategy to progressively optimize the model.
-Our model is built upon the open-source I2V model DynamiCrafter, which consists of a VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ for image compression, a UNet with spatial and temporal layers, and CLIP text and image encoders.
-We employ a dual-stream injection approach where the reference image is encoded by CLIP image encoder and injected into UNet through image cross-attention.
-Simultaneously, the first frame of the point cloud rendering is replaced with the reference image, which is then encoded by the VAE encoder and concatenated with noise before being fed into the UNet.
-This design ensures seamless integration of spatial and temporal information for high-quality video generation.
-3.2 Model Architecture
+The model builds on Video Diffusion Model (VDM) and consists of three main components: the Image2Cloud reconstructs 3D point cloud from a single reference image and generates point cloud renderings along a user-defined camera trajectory; the ObjMotionNet injects object dynamics into the UNet by encoding sparse trajectories into multi-scale motion features; the Spatial Triple-Attention Transformer integrates image, text, and lighting information via parallel cross-attention modules.
+The model enables I2V generation conditioned on arbitrary combinations of camera motion, object motion, and lighting direction.
+VidCRAFT3 is, to the best of our knowledge, the first image-to-video (I2V) diffusion framework that enables both independent and joint control of camera motion, object motion, and lighting direction, as illustrated in Fig.
+2.
+Given a reference image $I_{\text{ref}}\in\mathbb{R}^{H\times W\times 3}$ and a text prompt, the model generates a video $I=\{I_{f}\}_{f=1}^{F}$ with $I_{1}=I_{\text{ref}}$.
+The video frames will be generated according to the text description and all the control signals provided.
+Control signals can be used singly or in combination, namely camera trajectory $E=\{E_{f}\}_{f=1}^{F}$ with $E_{f}\in SE(3)$, sparse object trajectories $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\}$ for up to $N$ objects, and lighting direction $L_{\text{ref}}\in\mathbb{R}^{3}$.
+The model builds on the I2V model DynamiCrafter [98] (Sec.
+III-C).
+Appearance is preserved by encoding $I_{\text{ref}}$ with the CLIP image encoder and injecting the embeddings through image cross-attention.
+Precise camera motion control is achieved by reconstructing a 3D point cloud from $I_{\text{ref}}$ with Image2Cloud and rendering along the camera trajectory $E$ to obtain geometry-aware renderings $R=\{R_{f}\}_{f=1}^{F}$.
+These renderings are encoded by VAE and concatenated with noise as the UNet input, which anchors global camera motion while allowing the video diffusion model (VDM) to refine appearance and temporal coherence.
+Object motion signal is provided by ObjMotionNet, which converts sparse trajectories into a dense smoothed motion tensor and injects multi-scale motion features into UNet encoder blocks so that structure is aligned without overconstraining details.
+Lighting direction is handled by a Spatial Triple Attention Transformer that integrates image, text, and lighting direction through parallel cross-attention layers.
+To address the scarcity of real-world datasets with fully joint annotations, we construct three specialized datasets (Sec.
+III-D) and adopt a three-stage training strategy (Sec.
+III-E) to progressively optimize the model.
+III-B Preliminary
+Video diffusion models (VDMs) represent a class of generative models that extend the principles of image diffusion to the domain of video generation.
+These models operate by defining a forward diffusion process that gradually transforms an initial video sample $x_{0}\sim p_{\text{data}}(x)$ into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ timesteps.
+The reverse process, parameterized by a denoising network $\epsilon_{\theta}(x_{t},t)$, learns to iteratively denoise the noisy latent representation $x_{t}$ to recover the original data $x_{0}$.
+The training objective is formulated as:
+$\min_{\theta}\mathbb{E}_{t,x,\epsilon\sim\mathcal{N}(0,I)}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2}\right],$
+where $\epsilon$ represents the ground truth noise, and $\theta$ denotes the learnable parameters of the network.
+Once trained, the model can generate high-quality videos by sampling from a random noise distribution $x_{T}$ and applying the learned denoising process iteratively.
+To address the computational challenges associated with high-dimensional video data, Latent Diffusion Models (LDMs) are often employed.
+In this framework, a video $x\in\mathbb{R}^{F\times 3\times H\times W}$ is first encoded into a lower-dimensional latent space $z=\mathcal{E}(x)$, where $z\in\mathbb{R}^{F\times C\times h\times w}$.
+The diffusion and denoising processes are then performed in this latent space, significantly reducing computational complexity.
+The denoising process is conditioned on additional inputs $\mathbf{c}$, such as text prompts or motion control signals, enabling the generation of videos that adhere to specific semantic or temporal constraints.
+The final video is reconstructed through a decoder $\hat{x}=\mathcal{D}(z)$.
+III-C Model Architecture
 Camera Motion Control via Point Cloud Rendering.
-To achieve precise camera motion control in I2V, VidCRAFT3 leverages the Image2Cloud module, which reconstructs a high-quality 3D point cloud of the scene from a single reference image.
+Directly injecting camera parameters into the UNet learns an implicit mapping of camera poses to dynamics [7, 6], often causing coarse control, weak 3D consistency, and poor generalization to unseen trajectories.
+Inspired by [41, 42, 44], VidCRAFT3 leverages the Image2Cloud module, which reconstructs a high-quality 3D point cloud of the scene from a single reference image to provide explicit 3D priors, and combines it with a VDM for photorealistic refinement.
 Specifically, we employ DUSt3R, an unconstrained stereo 3D reconstruction model, to generate a 3D point cloud.
 Given a reference image $I_{\text{ref}}$, DUSt3R performs monocular or binocular reconstruction via point regression, followed by global alignment to ensure multi-view consistency: $\mathcal{P}=\text{DUSt3R}(I_{\text{ref}})$.
-The reconstructed point cloud provides explicit 3D geometry, enabling accurate rendering of the scene from arbitrary camera motion trajectory $E=\{E_{1},E_{2},\ldots,E_{F}\}$, where $F$ is the length of trajectory.
-Given a user-defined trajectory $E$, the point cloud rendering at frame $t$ is computed as $R_{t}=\pi(\mathcal{P},E_{t})$, where $\pi(\cdot)$ is the differentiable rendering function.
-However, due to the limitations of point cloud representation and the sparse 3D cues from a single image, the point cloud renderings may exhibit artifacts such as missing regions, occlusions, and geometric distortions.
-To address this, VidCRAFT3 integrates point cloud renderings as a input of VDM, which refines the coarse renderings to generate high-quality and temporally consistent video frames.
-This combination of explicit 3D geometry and generative refinement ensures both accurate camera control and realistic video synthesis.
+The reconstructed point cloud provides explicit 3D geometry, enabling accurate rendering of the scene from arbitrary camera trajectory.
+Given a user-defined camera trajectory $E=\{E_{f}\}_{f=1}^{F}$ with $E_{f}\in SE(3)$, the point cloud rendering at frame $t$ is computed as $R_{f}=\pi(\mathcal{P},E_{f})$, where $\pi(\cdot)$ is the differentiable rendering function.
+We replace the first point cloud rendering with the reference image, i.e., $R_{1}=I_{\text{ref}}$, which enforces an exact first-frame match and reduces drift from point-cloud noise and camera pose.
+Concretely, we use the per-frame rendering as the camera-conditioning input, i.e., $c_{\text{cam}}=\{R_{f}\}_{f=1}^{F}$.
+Following DynamiCrafter, we encode the point cloud rendering with the VAE, sample noise $\epsilon_{t}\!\sim\!\mathcal{N}(0,I)$, and concatenate it channel-wise to form the model input in latent space.
+Due to the limitations of point cloud representation and the sparse 3D cues from a single image, the point cloud renderings may exhibit artifacts such as missing regions, occlusions, and geometric distortions.
+To address this, VidCRAFT3 integrates point cloud renderings as an input to the VDM, which refines the coarse renderings to generate high-quality and temporally consistent video frames.
+This combination of explicit 3D geometry and VDM ensures both accurate camera control and realistic video synthesis.
 Object Motion Control through Trajectory Learning.
-Object motion in VidCRAFT3 is controlled through sparse spatial trajectories.
-For up to $N$ objects in a $F$-frames video, each trajectory is defined as a sequence of 2D pixel coordinates $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\mid n\in[1,N],\,f\in[1,F]\}$, where $\mathbf{s}_{n}^{f}$ denotes the position of the $n$-th object in frame $f$.
-To model motion dynamics, we compute inter-frame optical flow vectors between consecutive frames.
-For each trajectory point $\mathbf{s}_{n}^{f}$, the displacement vector $\mathbf{v}_{n}^{f}$ is calculated as $\mathbf{v}_{n}^{f}=\mathbf{s}_{n}^{f+1}-\mathbf{s}_{n}^{f}=(x_{n}^{f+1}-x_{n}^{f},\,y_{n}^{f+1}-y_{n}^{f}).$ These sparse motion vectors are then projected into a per-frame optical flow map $\mathcal{V}^{f}\in\mathbb{R}^{H\times W\times 2}$, where $H$ and $W$ denote spatial dimensions.
+Object motion in VidCRAFT3 is controlled through sparse object trajectories that are drawn by the user on the reference image.
+For up to $N$ objects in an $F$-frames video, each trajectory is defined as a sequence of 2D pixel coordinates $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\}$, with $n\in\{1,\ldots,N\}$ and $f\in\{1,\ldots,F\}$.
+$\mathbf{s}_{n}^{f}$ denotes the position of the $n$-th object in frame $f$.
+To model motion dynamics, we compute inter-frame optical flow vectors.
+For each trajectory point $\mathbf{s}_{n}^{f}$, the displacement vector $\mathbf{v}_{n}^{f}$ is calculated as $\mathbf{v}_{n}^{f}=\mathbf{s}_{n}^{f+1}-\mathbf{s}_{n}^{f}=(x_{n}^{f+1}-x_{n}^{f},\,y_{n}^{f+1}-y_{n}^{f}),f\in\{1,\ldots,F-1\}.$ These sparse motion vectors are then projected onto a per-frame optical flow map $\mathcal{V}^{f}\in\mathbb{R}^{H\times W\times 2}$.
 The mapping is formalized as
@@
 The full spatiotemporal flow tensor $\mathcal{V}\in\mathbb{R}^{F\times H\times W\times 2}$ is subsequently processed through Gaussian smoothing to obtain a dense motion representation $\tilde{\mathcal{V}}$.
-The ObjMotionNet is a neural network composed of multiple convolutional layers and downsampling operations, designed to extract multi-scale motion features from $\tilde{\mathcal{V}}$.
+The ObjMotionNet is a neural network composed of multiple convolutional layers and downsampling operations, designed to extract multi-scale motion features $c_{\text{obj}}$.
 Inspired by T2I-Adapter [99], ObjMotionNet injects multi-scale motion features exclusively into the UNet encoder.
 This balances precise motion control with video quality, as the encoder handles structure while the decoder refines details, ensuring accurate guidance without compromising output quality.
-Figure 3: Construction pipeline of the Object Motion Control Dataset.
 Lighting Direction Control with Spatial Triple-Attention Transformer.
-The lighting direction is represented as a per-frame 3D vector $L=(l_{x},l_{y},l_{z})$, describing the orientation of light source in Cartesian coordinates.
+In VidCRAFT3, the lighting direction is specified by the user as a unit vector $L_{\text{ref}}=(l_{x},l_{y},l_{z})$ defined in the reference view’s camera coordinate system.
+To derive lighting directions for other viewpoints along the camera trajectory, we transform $L$ using the extrinsic parameters.
+First, we promote $L$ to a homogeneous coordinate $\tilde{\mathbf{p}}_{\mathrm{ref}}=[L^{\top},1]^{\top}$ and use the inverse of the reference extrinsic $E_{\mathrm{ref}}^{-1}=E_{1}^{-1}$ to obtain its world-space position: $\tilde{\mathbf{p}}_{\mathrm{world}}=E_{\mathrm{ref}}^{-1}\,\tilde{\mathbf{p}}_{\mathrm{ref}}$.
+For each subsequent frame, world-to-camera extrinsic $E_{f}$ maps $\tilde{\mathbf{p}}_{\mathrm{world}}$ into that camera’s coordinate system, and we normalize the resulting three-vector to obtain a unit lighting direction:
+$\tilde{\mathbf{p}}_{f}=E_{f}\,\tilde{\mathbf{p}}_{\mathrm{world}},\quad L_{f}=\frac{[\tilde{\mathbf{p}}_{f}]_{1:3}}{\|[\tilde{\mathbf{p}}_{f}]_{1:3}\|}.$
+The sequence $L=\{L_{f}\}_{f=1}^{F}$ thus forms the per-frame lighting directions used in our model.
 To effectively encode this directional information into a high-dimensional feature space, we employ Spherical Harmonic (SH) Encoding.
 SH encoding captures the angular characteristics of the lighting using basis functions up to degree 4, resulting in 16 coefficients.
-The resulting SH-encoded vector $L_{\text{SH}}\in\mathbb{R}^{16}$ is projected into the feature space of the UNet using a multi-layer perceptron (MLP).
-$\mathbf{E}_{\text{light}}=\text{MLP}(L_{\text{SH}})$, where $\mathbf{E}_{\text{light}}$ is a lighting embedding aligned with the dimensionality of text embedding.
+The resulting SH-encoded vector $L_{\text{SH}}\in\mathbb{R}^{F\times 16}$ is projected into the feature space of the UNet using a multi-layer perceptron (MLP).
+$c_{\text{light}}=\text{MLP}(L_{\text{SH}})$, where $c_{\text{light}}$ is a lighting embedding aligned with the dimensionality of the text embedding.
 To incorporate the lighting embedding into the UNet, we propose the Spatial Triple-Attention Transformer, which integrates three parallel attention modules: image cross-attention, text cross-attention, and lighting cross-attention.
@@
 $\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (2)
-where $Q$ (Query) comes from the self-attention output of the UNet, $K$ (Key) and $V$ (Value) are derived from $\mathbf{E}_{\text{light}}$.
-The outputs of three cross-attention are summed to produce a fused feature representation $\mathbf{O}=\mathbf{O}_{\text{image}}+\mathbf{O}_{\text{text}}+\mathbf{O}_{\text{light}}$ where $\mathbf{O}_{\text{image}}$, $\mathbf{O}_{\text{text}}$, and $\mathbf{O}_{\text{light}}$ are the outputs of the image, text, and lighting cross-attention modules, respectively.
+where $Q$ (Query) comes from the self-attention output of the UNet, $K$ (Key) and $V$ (Value) are derived from $c_{\text{light}}$.
+The outputs of the three cross-attention modules are summed to produce a fused feature representation $\mathbf{O}=\mathbf{O}_{\text{image}}+\mathbf{O}_{\text{text}}+\mathbf{O}_{\text{light}}$, where $\mathbf{O}_{\text{image}}$, $\mathbf{O}_{\text{text}}$, and $\mathbf{O}_{\text{light}}$ are the outputs of the image, text, and lighting cross-attention modules, respectively.
 This novel mechanism ensures the generated videos maintain consistency across lighting, text, and image conditions.
-Figure 4: Construction pipeline of the VLD Dataset
-3.3 Dataset Construction
-Due to the lack of datasets with annotations for camera motion trajectory, object motion trajectories, and lighting directions, we construct three specialized datasets.
+III-D Dataset Construction
+Due to the lack of datasets annotated with camera motion trajectories, object motion trajectories, and lighting directions, we construct three specialized datasets.
 All datasets consist of 25-frame video clips with a spatial resolution of $320\times 512$ pixels.
 Camera Motion Control Dataset.
-We constructe this dataset from RealEstate10K [75], creating 62,000 clips with smooth camera transitions.
-Using DUSt3R, we generated globally aligned point clouds and rendered the first frame’s point cloud along the ground-truth camera trajectory.
-Since RealEstate10K lacks captions, we uniformly sampled 4 frames per clip and generated descriptions using Qwen2-VL-7B-Instruct [53].
-Despite its focus on indoor scenes, the dataset provides diverse camera trajectories for fine-grained motion control.
+We construct this dataset from RealEstate10K [100], curating 62,000 clips with smooth camera trajectories.
+As shown in Fig.
+3, for each clip, we take the first frame as the reference image, use DUSt3R to reconstruct a globally aligned 3D point cloud, and render the point cloud along the ground-truth camera trajectory to produce geometry-aware renderings.
+Since RealEstate10K lacks captions, we uniformly sample 4 frames from each clip and use Qwen2-VL-7B-Instruct [101] to generate a clip-level caption.
+Although RealEstate10K is predominantly indoors, the collected clips cover diverse camera motions, providing rich supervision for fine-grained camera control.
+Figure 3: Construction pipeline of the Camera Motion Control Dataset.
+We reconstruct a 3D point cloud from the first frame with DUSt3R, render views along the ground-truth camera trajectory, and generate a clip-level caption with Qwen2-VL-7B-Instruc.
 Object Motion Control Dataset.
@@
 The dataset construction pipeline, illustrated in Fig.
-3, consists of the following five steps: (1) Clip Filtering: Clips with abrupt scene changes are removed using PySceneDetect, and sequences with temporal intervals (1–16 frames) are sampled.
+4, follows these five steps: (1) Clip Filtering: Clips with abrupt scene changes are removed using PySceneDetect [^1] https://github.com/Breakthrough/PySceneDetect, and sequences with temporal intervals (1–16 frames) are sampled.
 To retain only clips with significant motion, optical flow is computed via MemFlow [102], and the bottom 25% of clips with low motion scores are filtered out.
-(2) Dense Trajectory Generation: For each filtered clip, CoTrackerV3 [24] generates dense trajectories on a $16\times 16$ grid, capturing pixel-level motion dynamics across frames.
-(3) Trajectory Filtering by Length: Trajectories shorter than the average length are discarded to ensure meaningful object motion.
+(2) Grid-point Tracking: For each filtered clip, we place a $16\times 16$ grid on the first frame and track each grid point across the clip using CoTrackerV3 [103].
+(3) Dense Trajectory Sampling: For each trajectory, the average per-frame displacement is computed and normalized by the image diagonal.
+A clip is discarded as camera-dominated if at least 60% of trajectories exceed 3% of the diagonal; otherwise, it is retained as object-motion-only.
+Within retained clips, trajectories shorter than the clip-level average length are removed to ensure meaningful object motion.
 (4) Sparse Trajectory Sampling: From the filtered dense trajectories, 1–8 sparse trajectories are sampled with probability proportional to their length.
@@
 Finally, a Gaussian filter is applied to smooth the sparse trajectory matrix, ensuring stable training.
+Figure 4: Pipeline for the Object Motion Control Dataset: clip filtering, grid-point tracking, dense trajectory sampling, sparse trajectory sampling, and optical flow smoothing.
 Lighting Direction Control Dataset.
-Since it is challenging to control a camera to follow the same trajectory under varying lighting conditions in real-world scenarios, collecting such datasets is both extremely difficult and costly.
-To address this, we introduce the VideoLightingDirection (VLD) Dataset, a novel synthetic dataset designed for modeling complex light-object interactions.
-The VLD Dataset comprises 57,600 synthetic videos rendered using Blender, featuring 3,600 realistic scenes, each with 16 videos captured under distinct lighting conditions.
+Collecting videos that follow the same camera trajectory under different lighting conditions is impractical in real-world settings, making such data extremely difficult and costly to obtain.
+We introduce VideoLightingDirection (VLD), a synthetic dataset of 57,600 Blender-rendered videos across 3,600 static scenes, each providing 16 distinct lighting directions under an identical camera trajectory.
 As illustrated in Fig.
-4, the dataset construction pipeline consists of the following four steps: (1) Scene Creation: We designed two scene types: Haven scenes and BOP scenes.
-Haven scenes places 3D models from Poly Haven at the center of HDR environments, while BOP scenes randomly places BOP models in a six-plane textured room to simulate indirect lighting.
-Models and environments can be randomly combined for diversity.
-See supplementary materials for BOP scenes examples.
-(2) Camera Trajectory Sampling: We sample the starting point of the camera trajectory on a spherical region (radius: 0.7–1.3 meters) centered around the 3D models.
-A smooth trajectory is then randomly generated around the model, ensuring that the camera always facing the center of the model.
+5, the construction pipeline comprises four steps: 1.
+Scene Creation.
+To better emulate real-world lighting scenarios, we design two complementary scene types: Haven and BOP.
+Haven uses Poly Haven HDR environment lighting together with a directional spotlight and places 3D models at the center of the HDR environment.
+BOP disables environment lighting, uses a single spotlight as the only light source, and randomly places BOP models in a six-plane textured room to simulate indirect lighting.
+Models and environments are randomly combined to increase diversity.
+(2) Camera Trajectory Sampling: We sample the starting position of the camera trajectory on a spherical shell centered at the 3D models with radius $r\in[0.7,1.3]$ meters.
+A smooth trajectory is then randomly generated around the model, ensuring that the camera always faces the center of the model.
 (3) Lighting Direction Sampling: To enhance the lighting effect, we reduce the HDR environment intensity by 40%.
-We uniformly sample 16 points on a hemisphere centered on the model, with the base surface normal aligned with the camera’s viewing direction.
-Each sampled point serves as the position of a 2kW spotlight (radius: 1), oriented toward the model’s center.
-The light source position vectors are mapped to all camera coordinate systems and normalized to compute the lighting direction for each view.
+We uniformly sample 16 points on a hemisphere centered on the model, with the base surface normal aligned with the reference camera viewing direction.
+Each sampled point serves as the position of a 2kW spotlight (radius=1), oriented toward the model’s center.
+The lighting direction vectors are projected to the corresponding camera coordinate and normalized to obtain per-frame lighting direction labels (Sec.
+III-C).
 (4) Rendering and Annotation: Using Blender Cycles, we render each scene under 16 different lighting conditions while maintaining a consistent camera trajectory.
-Each frame is annotated with the corresponding camera pose and lighting direction, providing precise control signals for training.
-3.4 Training Strategy
-To enable simultaneous control over camera motion, object motion, and lighting direction, we adopt a three-stage progressive training strategy, designed to address the absence of datasets annotated with all three control signals.
+Each frame is annotated with the camera trajectory and lighting direction, providing precise supervision signals for training.
+Fig.
+6 shows samples from the VideoLightingDirection (VLD) dataset.
+We include two types of scenes: (a) Haven scenes, in which 3D models from Poly Haven are placed centrally within HDR environments; and (b) BOP scenes, featuring randomly positioned BOP models in six-plane textured rooms to simulate indirect lighting.
+We display two lighting directions per scene, while the dataset actually provides 16 distinct lighting directions under an identical camera trajectory.
+This design highlights how lighting direction affects shading, reflections, and visual coherence, offering valuable data for training models with precise lighting annotations.
+Figure 5: Construction pipeline for the VideoLightingDirection Dataset.
+A 3D model is randomly paired with an HDR environment to create a static scene; a smooth camera trajectory and a single spotlight are then sampled around the 3D model and rendered in Blender to produce a video clip with the lighting direction annotation.
+III-E Training
+We train VidCRAFT3 conditioned on text, a reference image, camera motion, object motion, and lighting direction, using a standard denoising diffusion objective.
+Since no single dataset provides all three control annotations jointly, we adopt a three-stage progressive training strategy.
+Training Objective.
+Let $z_{0}$ denote the clean video latent of a training sample $x$, and let $c=\{c_{\text{img}},c_{\text{txt}},c_{\text{cam}},c_{\text{obj}},c_{\text{light}}\}$ denote the control signals.
+Here, $c_{\text{txt}},c_{\text{img}},c_{\text{cam}},c_{\text{obj}}$, and $c_{\text{light}}$ refer to text prompt, reference-image latent, point cloud renderings, multi-scale motion features, and lighting embedding, respectively.
+We train the noise estimator $\epsilon_{\theta}$ to reverse the diffusion process:
+$\min_{\theta}\;\mathbb{E}_{z_{0},c,t,\epsilon\sim\mathcal{N}(0,I)}\!\left[\big\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\big\|_{2}^{2}\right],$ (3)
+where $\epsilon$ denotes Gaussian noise, $t\!\in\!\{0,\ldots,T\}$ is the diffusion timestep, and $z_{t}$ is the noised latent at step $t$.
+During training, we apply classifier-free dropout: with probability $p_{\text{uncond}}$, we discard one randomly chosen conditioning branch or discard all; if $c_{\text{cam}}$ is discarded, it is replaced by the reference-image latent repeated $F$ times to match the temporal length (following DynamiCrafter), whereas other dropped conditionings are replaced with zero tensors of matching shape.
+Figure 6: Illustrations of samples from the proposed VideoLightingDirection (VLD) Dataset, featuring synthetic scenes designed to model complex light-object interactions.
+(a) Haven scenes, where 3D models from Poly Haven are placed at the center of HDR environments.
+(b) BOP scenes, with BOP models randomly positioned within six-plane textured rooms to simulate indirect lighting.
+Each subset includes video frames captured under two distinct lighting conditions for visualization, maintaining consistent camera trajectories.
 Stage 1: Camera Motion Control Training.
 The training begins by initializing the model with DynamiCrafter pre-trained weights.
-The model is then fine-tuned on camera motion control dataset for 40,000 iterations, optimizing the entire UNet to align VDM with camera motion.
-This stage establishes robust 3D scene understanding by integrating point cloud renderings, ensuring precise global camera movements while maintaining temporal consistency.
+The model is then fine-tuned on camera motion control dataset for 40,000 iterations.
+We optimize only the UNet while freezing the other modules to align the VDM with camera motion.
+This stage establishes robust 3D scene understanding by integrating point cloud renderings and aligns the VDM with precise global camera motion, while maintaining temporal consistency.
 Stage 2: Dense Object Trajectories and Lighting Mixed Fine-tuning.
@@
 This hybrid dataset enhances the ability of model to learn joint control over these three conditions.
+We use dense object trajectories obtained from the “Trajectory Filtering by Length” step of the Object Motion Control Dataset pipeline; these dense trajectories are then processed by the same Optical Flow Smoothing procedure used for sparse trajectories to yield stable dense motion fields.
 Dense object trajectories are incorporated to provide rich motion details, accelerating model convergence.
@@
 The same hybrid dataset from Stage 2 is reused, but dense trajectories are replaced with sparse trajectories to simulate real-world user interactions.
-Fine-tuning follows the same parameters as Stage 2, with 20,000 iterations.
+Fine-tuning the same parameters as Stage 2, with 20,000 iterations.
 This stage forces the model to infer complex motion patterns from limited trajectory data, and the progressive shift from dense to sparse supervision enhances its ability to generalize to practical scenarios.
-By progressively learning camera motion, object motion, and lighting direction controls across stages while strategically freezing layers to retain prior knowledge, the training strategy ensures that VidCRAFT3 achieves fine-grained, synergistic control over all three elements, even in the absence of fully annotated multi-task training data.
-4 Experiments
-Figure 5: Qualitative comparisons with SOTA methods on RealEstate10K.
-Figure 6: Qualitative comparisons with SOTA methods on WebVid-10M.
-4.1 Experimental Setup
+By progressively learning camera motion, object motion, and lighting direction controls across stages while strategically freezing layers to retain prior knowledge, the training strategy yields fine-grained, synergistic control over all three elements without requiring fully annotated multi-task data.
+III-F Inference
+During inference, given a reference image $I_{\text{ref}}$ and a text prompt, users may optionally input one control or multiple controls, including camera trajectory $E$, sparse object trajectories $\mathcal{T}$, and a lighting direction $L_{\text{ref}}$.
+We reconstruct a 3D point cloud from $I_{\text{ref}}$ using Image2Cloud and render it along the camera trajectory to obtain geometry-aware renderings $R$ for camera motion.
+The sparse object trajectories $\mathcal{T}$ are encoded by ObjMotionNet into multi-scale motion features.
+The reference lighting direction $\mathbf{L}_{\text{ref}}$ is SH encoded and projected via an MLP, then fused through the Spatial Triple-Attention Transformer.
+In addition, text and image embeddings modulate the U-Net via the same transformer.
+Consistent with training, any missing condition is replaced by its “null” signal: when camera control is absent, a stationary camera is equivalent to rendering the reference view $F$ times, and for efficiency we directly repeat the reference image latent across $F$ frames; when object or lighting control is absent, we input zero tensors with matching shapes.
+This enables each control to operate alone or in any combination.
+We sample with DDIM and classifier-free guidance to generate video.
+IV Experiments
+Figure 7: Qualitative comparison of camera motion control on RealEstate10K.
+The left column shows target camera trajectory visualizations; rows show ground-truth sequences (GT), CameraCtrl[7], MotionCtrl[6], CamI2V[33], and Ours.
+VidCRAFT3 follows the camera trajectory more faithfully, preserving layout and parallax with fewer artifacts.
+Figure 8: Qualitative comparison of object motion control on WebVid-10M.
+The left column shows target object trajectories visualizations; rows show ground-truth sequences (GT), Image Conductor [55], Motion-I2V [58], and Ours.
+VidCRAFT3 maintains identity and background while adhering to specified object trajectories, producing fewer artifacts and jitter across challenging scenes.
+IV-A Experimental Setup
 Implementation Details.
@@
 During training and inference, video clips are processed at $320\times 512$ resolution with 25 frames.
-To optimize the model, we employ the Adam optimizer with a learning rate of $1\times 10^{-5}$ and a batch size of 96.
-The training is conducted on 8 NVIDIA H100 GPUs.
-For inference, we adopt the DDIM sampler and classifier-free guidance to enhance generation quality and multimodal consistency.
-During inference, the model averagely uses 20 GB of GPU memory and takes 42 seconds.
-[…]
-CameraCtrl 97.99 96.11 29.41 4.19
-MotionCtrl 103.82 188.93 30.18 4.23
-CamI2V 98.54 85.03 30.37 4.24
-Ours 75.62 49.77 32.32 4.07
-Table 1: Quantitative comparison with SOTA methods on RealEstate10K.
+We optimize the model using Adam with a learning rate of $1\times 10^{-5}$ and a global batch size of 96, and adopt classifier-free training with an unconditional drop probability of $0.05$.
+Training is conducted on 8 NVIDIA H100 GPUs.
+For inference, we use a DDIM sampler and classifier-free guidance with a guidance scale of $7.5$.
+On average, inference uses about 20 GB of GPU memory and takes roughly 42 seconds per sample.
 Evaluation Datasets.
@@
 For object motion control, we sample 1,000 samples from the WebVid-10M test set.
-For lighting direction control, we sample 1,000 samples from the VLD dataset, covering a wide range of lighting directions.
-To provide a broader evaluation, we create a generalized test set consisting of 100 videos sourced from copyright-free websites like Pixabay and Pexels, as well as videos generated by T2V models.
+For lighting direction control, we evaluate on 1,000 samples from VLD, stratified as 500 Haven-type scenes and 500 BOP-type scenes, covering a wide range of lighting directions.
+To provide a broader evaluation, we create a generalized test set consisting of 100 videos sourced from copyright-free websites such as Pixabay and Pexels, as well as videos generated by T2V models.
 This dataset spans categories such as human activities, animals, vehicles, indoor scenes, artworks, natural landscapes, and AI-generated images.
@@
 We evaluate VidCRAFT3 across three dimensions: (1) Video Quality: FID for assessing visual fidelity, FVD for measuring temporal coherence, and CLIPSIM for detecting semantic alignment; (2) Motion Control Performance: Based on camera poses estimated by DUSt3R and object trajectories extracted by CoTrackerV3, motion control performance is quantified using CamMC and ObjMC metrics [6], which measure the Euclidean distance between predicted and ground-truth values; (3) Lighting Control Effectiveness: Evaluated through LPIPS, SSIM, and PSNR, comparing generated frames with ground-truth images to assess perceptual quality and structural fidelity.
-4.2 Comparisons with State-of-the-Art Methods
-Due to the lack of open-source I2V methods capable of simultaneously controlling camera and object motion, we evaluate VidCRAFT3 separately for camera motion control and object motion control, comparing it against SOTA methods in each domain.
+Figure 9: Qualitative results of the same reference image under different camera motion, object motion and lighting direction.
+For each subfigure, we vary only one control while keeping the other two fixed: (a) camera motion, (b) object motion, and (c) lighting direction.
+The leftmost panels visualize the target control (camera trajectory / object trajectory / lighting direction).
+VidCRAFT3 faithfully follows the control signal while preserving content fidelity.”
+IV-B Main Results
+VidCRAFT3 demonstrates strong controllability and visual fidelity across camera motion, object motion, and lighting direction, while preserving temporal and semantic consistency.
+As no open-source I2V method supports simultaneous camera–object control, we evaluate VidCRAFT3 separately on camera motion and object motion against state-of-the-art baselines, then present qualitative joint control results.
+TABLE I: Quantitative comparison of camera motion control on RealEstate10K. VidCRAFT3 shows stronger camera adherence with improved visual fidelity, temporal coherence, and semantic alignment.
 […]
-Image Conductor 150.26 242.01 29.69 12.96
-Motion-I2V 128.35 171.35 30.92 3.96
+CameraCtrl [7] 97.99 96.11 29.41 4.19
+MotionCtrl [6] 103.82 188.93 30.18 4.23
+CamI2V [33] 98.54 85.03 30.37 4.24
+Ours 75.62 49.77 32.32 4.07
+TABLE II: Quantitative comparison of object motion control on WebVid-10M.
+VidCRAFT3 shows stronger object adherence and identity preservation, with improved visual fidelity, temporal coherence, and semantic alignment
+[…]
+Image Conductor [55] 150.26 242.01 29.69 12.96
+Motion-I2V [58] 128.35 171.35 30.92 3.96
 Ours 87.12 120.65 32.99 3.51
-Table 2: Quantitative comparison with SOTA methods on WebVid-10M.
-Camera Motion Control.
-VidCRAFT3 demonstrates superior performance in camera motion control compared to state-of-the-art (SOTA) methods like CameraCtrl, CamI2V, and MotionCtrl on the RealEstate 10K dataset.
-As shown in Table 1 and Fig.
-5, VidCRAFT3 excels in camera motion control, as evidenced by its CamMC score of 4.07, which is lower than CameraCtrl (4.19), CamI2V (4.24), and MotionCtrl (4.23).
-VidCRAFT3 also achieves better results in other key metrics indicate that VidCRAFT3 not only controls camera motion more accurately but also generates videos with higher visual quality, temporal coherence, and semantic alignment.
-Qualitatively, VidCRAFT3 generates smoother and more realistic camera movements with fewer artifacts, particularly in complex scenes.
-This success stems from the Image2Cloud module, which enable precise 3D scene reconstruction and fine-grained camera control.
-Object Motion Control.
-VidCRAFT3 demonstrates exceptional performance in object motion control, as highlighted in Table 2.
-The model achieves an ObjMC score of 3.51, which is lower than Image Conductor (12.96) and Motion-I2V (3.96).
-This indicates that VidCRAFT3 more accurately aligns the generated object trajectories with the ground truth, resulting in more realistic and faithful object motion.
-Additionally, VidCRAFT3 outperforms in other key metrics, showcasing superior visual quality, temporal coherence, and semantic alignment.
+IV-B1 Camera Motion Control
+VidCRAFT3 achieves accurate and stable camera motion control on RealEstate10K.
+As shown in Fig.
+9(a), under different complex camera trajectory controls, the model accurately generates videos with the specified camera motion while preserving scene content.
+Quantitatively, Table I reports the best results across all metrics: CamMC 4.07 (better than the best baseline, 4.19), FID 75.62, FVD 49.77, and CLIPSIM 32.32, outperforming CameraCtrl, CamI2V, and MotionCtrl.
+These improvements indicate more precise camera control, along with better visual quality, temporal coherence, and semantic alignment.
+Qualitatively, Fig.
+7 shows smoother and more realistic camera motion with fewer artifacts, particularly in complex scenes.
+We attribute these improvements to the explicit 3D priors provided by Image2Cloud, which anchor camera motion to a consistent scene geometry and significantly reduce drift during large viewpoint changes.
+IV-B2 Object Motion Control
+VidCRAFT3 demonstrates exceptional performance in object motion control on WebVid-10M.
+As shown in Fig.
+9(b), under different complex object trajectories controls, the model can accurately control the object motion in the video.
+Quantitatively, Table II reports that the model achieves an ObjMC score of 3.51, which is lower than Image Conductor (12.96) and Motion-I2V (3.96).
+This indicates that VidCRAFT3 achieves closer alignment between generated and target trajectories, resulting in more realistic and faithful object motion.
+Additionally, VidCRAFT3 outperforms other methods in other metrics, showcasing superior visual quality, temporal coherence, and semantic alignment.
 Qualitatively, as shown in Fig.
-6, VidCRAFT3 generates more realistic and consistent object movements compared to Image Conductor and Motion-I2V.
+8, VidCRAFT3 generates more realistic and consistent object movements compared to Image Conductor and Motion-I2V.
 The model effectively captures the dynamics of object motion, ensuring smooth transitions and natural interactions within the scene.
 These results highlight VidCRAFT3’s robust object motion control, driven by its advanced ObjMotionNet, which effectively captures and controls complex motion patterns.
+TABLE III: Quantitative results for lighting direction control on VLD across 1,000 clips, stratified by scene type: 500 Haven-type and 500 BOP-type.
 […]
-CameraCtrl 3.4% 5.4% 4.7%
-MotionCtrl 5.7% 7.3% 6.5%
-CamI2V 5.6% 6.8% 5.2%
+Haven-type 105.70 123.65 20.45 17.79 0.67 0.13 5.21
+BOP-type 95.96 111.73 26.95 21.19 0.81 0.09 4.79
+Overall 100.83 117.69 23.70 19.49 0.74 0.11 5.00
+TABLE IV: User study. Results demonstrating our method’s superior performance in camera and object motion control compared to baseline approaches across all metrics.
+[…]
+CameraCtrl [7] 3.4% 5.4% 4.7%
+MotionCtrl [6] 5.7% 7.3% 6.5%
+CamI2V [33] 5.6% 6.8% 5.2%
 Ours 85.3% 80.5% 83.6%
 […]
-Image Conductor 10.3% 11.9% 11.2%
-Motion-I2V 14.8% 10.9% 13.2%
+Image Conductor [55] 10.3% 11.9% 11.2%
+Motion-I2V [58] 14.8% 10.9% 13.2%
 Ours 74.9% 77.2% 75.6%
-Table 3: User study. Results demonstrating our method’s superior performance in camera and object motion control compared to baseline approaches across all metrics.
+Figure 10: Qualitative results demonstrating simultaneous control over camera motion and object motion.
+For each block, we show the specified camera trajectory and object trajectories as controls, followed by frames from the ground-truth video (GT, top) and from VidCRAFT3 (Ours, bottom).
+IV-B3 Lighting Direction Control
+VidCRAFT3 enables accurate and realistic lighting direction control for both synthetic and real reference images.
+As shown in Fig.
+9(c), with camera and object motion fixed, supplying a sequence of lighting directions produces videos whose illumination follows the control signal while preserving appearance.
+Notably, although training samples contain only static lighting directions, the model produces coherent responses to time-varying direction sequences at test time, demonstrating the combined generalization of direction-level conditioning and temporal modeling.
+Prior methods largely target portraits and rely on HDR environment maps rather than interactive, direction-level control in general scenes.
+Their evaluation protocols are also incompatible with our VLD benchmark.
+Accordingly, we report results on VLD and verify design choices via ablations (Table VIII, Table VII).
+Table III presents metrics on the Haven and BOP scene types: VidCRAFT3 performs comparably on both, with slightly better results on BOP.
+This suggests that a single spotlight is easier to disentangle than environment light + spotlight, while the small gap indicates that our model exhibits scene-agnostic domain generalization.
+IV-B4 Joint Control
+VidCRAFT3 delivers accurate and coherent joint control of camera motion, object motion, and lighting direction.
+We use the same inference setup as in the single-control setting and compose a subset of controls in a single forward pass.
+Qualitative results are shown in Fig.
+1, Fig.
+12, Fig.
+11, and Fig.
+10.
+For camera + lighting (Fig.
+12, Fig.
+11), the model follows the prescribed camera trajectory while maintaining consistent shading, highlights, and cast shadows across viewpoints, indicating temporally stable lighting.
+For camera + object (Fig.
+10), the model preserves the global camera trajectory and adheres to sparse object trajectories, retaining identity and shape with parallax that matches the camera motion.
+With all three controls (Fig.
+1), object motion aligns with target trajectories, lighting changes remain coherent with viewpoint transitions, and the scene layout stays stable under large viewpoint changes.
+These results demonstrate clean compositionality without interference among controls and sustained temporal consistency across frames.
+Figure 11: Qualitative results demonstrating simultaneous control over camera motion and lighting direction.
+For each row, we show a reference image, the specified camera trajectory and lighting direction controls, and the relighting video frames generated by VidCRAFT3.
+Figure 12: Qualitative results demonstrating simultaneous control over camera motion and lighting direction on VLD.
+The top block shows a Haven-type scene and the bottom block shows a BOP-type scene.
+Our generated videos (”Ours”) are compared against ground-truth sequences (”GT”).
+IV-C User Study
+To evaluate VidCRAFT3, we conducted a user study with 10 participants assessing 200 randomly selected results, comprising 100 camera motion samples (50 from copyright-free and 50 from the RealEstate10K test split) and 100 object motion samples (50 from copyright-free and 50 from the WebVid-10M test split), drawn from the outputs of our method and five baselines.
+The participants evaluated the results based on four metrics: Camera Motion Precision, Object Motion Precision, Visual Quality, and Overall Quality.
+As shown in Table IV, our method achieved over 80% in camera motion control and over 74% in object motion control across all metrics, demonstrating precise and visually appealing control, validating its effectiveness and robustness in real-world applications.
+IV-D Ablation Studies
 […]
-Text Cross-Atten 111.08 121.95 22.77 18.14 0.72 0.13 5.31
-Time Embed. 101.71 123.31 22.70 19.07 0.73 0.12 5.21
-Lighting Cross-Attn 100.83 117.69 23.70 19.49 0.74 0.11 5.00
-Table 4: Ablation of lighting embedding integration strategies on VLD.
-Text Cross-Atten method concatenates lighting embedding with text embedding and integrates them into the model through text cross-attention.
-Time Embed method adds lighting embedding to time embedding.
-The proposed Lighting Cross-Atten method, detailed in Sec.
-3.2, demonstrates enhanced performance in integrating lighting conditions.
-User Study.
-To evaluate our method, we conducted a user study with 10 participants assessing 200 randomly selected results from our method and five baselines.
-The participants evaluated the results based on four key metrics: Camera Motion Precision, Object Motion Precision, Visual Quality, and Overall Quality.
-As shown in Table 3, our method achieved over 80% in camera motion control and over 74% in object motion control across all metrics, demonstrating precise and visually appealing control, validating its effectiveness and robustness in real-world applications.
-4.3 Ablation Study
-Training Strategy.
-We compare training with Dense, Sparse, and Dense+Sparse trajectories.
-As shown in Table 5, Dense training (ObjMC: 4.39) provides rich motion information but struggles during inference due to sparse inputs.
-Sparse training (ObjMC: 4.05) improves inference alignment but lacks motion detail.
-The Dense+Sparse approach achieves the best results (ObjMC: 3.51), leveraging dense trajectories for robust learning and fine-tuning with sparse trajectories for better generalization.
-This hybrid strategy enhances object motion control and video quality, as shown by superior FID (87.12) and FVD (120.65) scores.
-Lighting Embedding Integration Strategies.
-We compare different lighting embedding integration strategies.
-As shown in Table 4, Lighting Cross-Atten achieves the best results, with an PSNR of 19.49, SSIM of 0.74, and LPIPS of 0.11, outperforming other methods.
-These findings confirm the effectiveness of Lighting Cross-Attention for precise lighting direction control.
-Representation of Lighting Direction.
-We compare Fourier Embedding [31] and SH Encoding for representation of lighting direction.
-As shown in Table 6, SH Encoding outperforms Fourier Embedding, achieving better PSNR (19.49), SSIM (0.74), and LPIPS (0.11).
-These results confirm SH Encoding’s effectiveness for precise lighting control in image-to-video generation.
-For more results, please refer to the supplementary materials.
+[…]
+Full three-Stage (S1 $\rightarrow$ S2 $\rightarrow$ S3)
+Stage 1 (camera) 3.98
+Stage 2 (object-dense + VLD) 4.19 4.39 18.21 0.60 0.20
+Stage 3 (object-sparse + VLD) 4.07 3.51 19.49 0.74 0.11
+No-S1 (S2 $\rightarrow$ S3)
+Stage 2 (dense + VLD) 6.52 4.35 18.47 0.68 0.21
+Stage 3 (sparse + VLD) 6.17 3.45 19.34 0.81 0.10
+[…]
 […]
@@
 Sparse 91.54 123.15 30.93 4.05
-Dense+Sparse 87.12 120.65 32.99 3.51
-Table 5: Ablation of training strategy on WebVid-10M.
+[…]
+Training Strategy.
+To validate the effectiveness of our three-stage training strategy, we conduct two experiments: (1) whether the progressive three-stage training strategy affects the capabilities learned in the previous stages (e.g., camera motion control), and (2) whether the camera prior learned in Stage 1 is necessary.
+As shown in Table V, in the full three-stage advancing from Stage 2 to Stage 3 improves object adherence and relighting quality while also improving camera adherence.
+In contrast, the No-S1 variant achieves similar object and relighting scores at Stage 3 but shows a pronounced drop in CamMC, indicating that Stage 1 provides a crucial camera and geometry prior.
+[…]
+Lighting Embedding Integration Strategies.
+We compare different lighting embedding integration strategies.
+(1) Text Cross-Attn method concatenates lighting embedding with text embedding and integrates them into the model through text cross-attention.
+(2) Time Embed method adds lighting embedding to time embedding.
+(3) The proposed Lighting Cross-Atten method, which introduces a dedicated lighting cross-attention branch inside the Spatial Triple-Attention Transformer.
+As shown in Table VII, Lighting Cross-Attn achieves the best results and outperforms the other methods across all metrics, indicating that explicit, decoupled lighting attention integrates lighting directions more effectively than routing them through text or time embedding.
+Representation of Lighting Direction.
+We compare Fourier Embedding [104] and SH Encoding for representation of lighting direction.
+The resulting lighting embeddings are fed to the Lighting Cross-Attn branch.
+As shown in Table VIII, SH Encoding outperforms Fourier Embedding across all metrics on VLD.
+We attribute these gains to the smooth, rotation-aware angular basis of SH, which provides a more stable and geometry-consistent signal for lighting than Fourier features.
+For more results, please refer to the supplementary materials.
+TABLE VII: Ablation of lighting embedding integration strategies on VLD. The proposed Lighting Cross-Attn achieves the best overall quality and lighting control.
+[…]
+Text Cross-Attn 111.08 121.95 22.77 18.14 0.72 0.13 5.31
+Time Embed. 101.71 123.31 22.70 19.07 0.73 0.12 5.21
+Lighting Cross-Attn 100.83 117.69 23.70 19.49 0.74 0.11 5.00
+TABLE VIII: Ablation of representation of lighting direction on VLD. SH Encoding surpasses Fourier Embedding across fidelity and alignment metrics.
 […]
@@
 SH Encoding 100.83 117.69 23.70 19.49 0.74 0.11 5.00
-Table 6: Ablation of representation of lighting direction on VLD.
-5 Conclusions
-We present VidCRAFT3, a novel framework for precise image-to-video generation with simultaneous control over camera motion, object motion, and lighting direction.
-It integrates Image2Cloud for 3D point cloud generation, ObjMotionNet for encoding sparse object trajectories, and a Spatial Triple-Attention Transformer for incorporating lighting direction.
-The VideoLightingDirection dataset and a three-stage training strategy address real-world annotation challenges.
-Experiments show that VidCRAFT3 outperforms state-of-the-art methods in control granularity and visual coherence, advancing video generation toward more realistic and versatile applications.
-Supplementary Material
-Appendix A Preliminary
-Image-to-video generation.
-Given a reference image $I_{\text{ref}}$ and some text description, image-to-video generation aims to generate a sequence of video frames $I=\{I_{1},I_{2},\ldots,I_{F}\}$ with $I_{1}=I$.
-In this work, we also allow the user to provide three forms of control signals: camera motion, object motion, and lighting direction.
-The video frames will be generated according to the text description and all the control signals provided.
-Specifically, the camera motion is represented by the movement of a set of camera extrinsics $E=\{E_{1},E_{2},\ldots,E_{F}\}$; the object motion is represented by a few sparse sequences of image pixels $\{s_{1}^{1},s_{1}^{2},s_{1}^{3},\ldots,s_{1}^{F}\},\{s_{1}^{2},s_{2}^{2},s_{2}^{3},\ldots,s_{2}^{F}\},\ldots$; the lighting direction is simply a unit 3d vector $L\in\boldsymbol{R}^{3}$.
-Video Diffusion model (VDMs) represent a class of generative models that extend the principles of image diffusion to the domain of video generation.
-These models operate by defining a forward diffusion process that gradually transforms an initial video sample $x_{0}\sim p_{\text{data}}(x)$ into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ timesteps.
-The reverse process, parameterized by a denoising network $\epsilon_{\theta}(x_{t},t)$, learns to iteratively denoise the noisy latent representation $x_{t}$ to recover the original data $x_{0}$.
-The training objective is formulated as:
-$\min_{\theta}\mathbb{E}_{t,x\sim p_{\text{data}},\epsilon\sim\mathcal{N}(0,I)}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2},$
-where $\epsilon$ represents the ground truth noise, and $\theta$ denotes the learnable parameters of the network.
-Once trained, the model can generate high-quality videos by sampling from a random noise distribution $x_{T}$ and applying the learned denoising process iteratively.
-To address the computational challenges associated with high-dimensional video data, Latent Diffusion Models (LDMs) are often employed.
-In this framework, a video $x\in\mathbb{R}^{F\times 3\times H\times W}$ is first encoded into a lower-dimensional latent space $\mathbf{z}=\mathcal{E}(x)$, where $\mathbf{z}\in\mathbb{R}^{F\times C\times h\times w}$.
-The diffusion and denoising processes are then performed in this latent space, significantly reducing computational complexity.
-The denoising process is conditioned on additional inputs $\mathbf{c}$, such as text prompts or motion control signals, enabling the generation of videos that adhere to specific semantic or temporal constraints.
-The final video is reconstructed through a decoder $\hat{x}=\mathcal{D}(\mathbf{z})$, ensuring both spatial and temporal coherence in the generated output.
-Appendix B VLD Dataset Samples
-Fig.
-7 illustrates representative samples from our proposed VideoLightingDirection (VLD) Dataset, a synthetic dataset specifically designed to model complex light-object interactions.
-The dataset comprises two types of scenes: (a) Haven scenes, in which 3D models from Poly Haven are placed centrally within HDR environments; and (b) BOP scenes, featuring randomly positioned BOP models in six-plane textured rooms to simulate indirect lighting.
-Each subset includes video frames captured under two distinct lighting conditions while maintaining a consistent camera trajectory.
-This design highlights how changes in lighting direction impact object shading, reflections, and overall visual coherence, providing valuable data for training and evaluating models requiring precise lighting direction annotations.
-Figure 7: Illustrations of samples from the proposed VideoLightingDirection (VLD) Dataset, featuring synthetic scenes designed to model complex light-object interactions.
-(a) Haven scenes, where 3D models from Poly Haven are placed at the center of HDR environments.
-(b) BOP scenes, with BOP models randomly positioned within six-plane textured rooms to simulate indirect lighting.
-Each subset includes video frames captured under two distinct lighting conditions, maintaining consistent camera trajectories.
-Appendix C Qualitative Results of Ablation Study
+V Conclusions
+In conclusion, we present VidCRAFT3, a unified and flexible framework for precise image-to-video generation that enables simultaneous control over camera motion, object motion, and lighting direction.
+VidCRAFT3 integrates three core components: Image2Cloud, which generates 3D point clouds from reference images; ObjMotionNet, which encodes sparse object trajectories using multi-scale optical flow features; and Spatial Triple-Attention Transformer, which incorporates lighting direction embeddings through parallel cross-attention modules.
+The introduction of the VideoLightingDirection dataset, combined with a three-stage training strategy, effectively addresses the challenges posed by the lack of annotated real-world datasets.
+Extensive experiments demonstrate that VidCRAFT3 produces high-quality video content, outperforming state-of-the-art methods in terms of control granularity and visual coherence.
+Our work advances the field of video generation by enabling enhanced control over multiple visual elements, thereby paving the way for more realistic and versatile applications.
+Acknowledgments
+This paper is supported by the Doubao Fund.
+Qualitative Results of Ablation Study
 In this section, we present qualitative results of the ablation studies described in the main text, providing visual insights into how different design choices impact the generated videos.
-Training Strategy.
-Fig.
-8 qualitatively compares videos generated using Dense, Sparse, and Dense+Sparse training strategies.
-Dense training results exhibit detailed object motion but struggle with precise alignment when tested with sparse trajectories.
-Sparse training achieves better alignment but at the cost of reduced motion detail.
-The Dense+Sparse approach effectively combines the strengths of both, delivering superior motion coherence and overall visual quality.
+[…]
 Lighting Embedding Integration Strategies.
 Fig.
-9 compares the generated videos of different lighting embedding integration strategies.
-Lighting Cross-Attention demonstrates clear superiority, providing precise control over lighting directions, resulting in more realistic shadows, reflections, and overall lighting effects compared to Text Cross-Attention and Time Embedding.
+14 qualitatively compares three lighting embedding integration strategies: Text Cross-Attention, Time Embedding, and Lighting Cross-Attention.
+Text Cross-Attention integrates lighting embeddings with text embeddings through cross-attention, resulting in less accurate lighting control and suboptimal lighting effects.
+Time Embedding improves upon this by incorporating lighting information into the time embedding, providing better consistency but still lacking in precision.
+Lighting Cross-Attention excels by directly integrating lighting embeddings through dedicated cross-attention, offering superior control over lighting directions and producing more realistic shadows, reflections, and overall lighting effects.
+This strategy aligns more closely with the ground truth (GT) and achieves better visual fidelity (cf.
+Table VII).
 Representation of Lighting Direction.
 Fig.
-10 presents a visual comparison between Fourier Embedding and SH Encoding for representing lighting direction.
-SH Encoding consistently produces more realistic and detailed lighting effects, including enhanced shading and reflections, further supporting the quantitative findings in the main text.
+15 qualitatively compares two methods for representing lighting direction: Fourier Embedding and SH Encoding.
+Fourier Embedding represents lighting direction using periodic basis functions but struggles with accurately capturing complex lighting effects.
+In contrast, SH Encoding leverages spherical harmonics to better model the angular properties of lighting, producing more realistic and detailed lighting effects, such as enhanced shading and reflections.
+These improvements result in outputs that align more closely with the ground truth (GT), as demonstrated in both visual and quantitative comparisons (cf.
+Table VIII).
 These qualitative analyses further validate our key design decisions, emphasizing their impact on generating precise and realistic image-to-video translations under controlled camera motion, object motion, and lighting direction.
-Appendix D Additional Qualitative Results
-This section provides additional qualitative comparisons illustrating VidCRAFT3’s capabilities across multiple control dimensions.
-Camera Motion + Lighting Direction Control.
-Fig.
-11 demonstrates VidCRAFT3’s effectiveness in simultaneously controlling camera motion and lighting direction, showing clear advantages in lighting realism and trajectory accuracy compared to original sequences.
-Camera Motion + Object Motion Control.
-Fig.
-12 presents examples of simultaneous control over camera and object motions.
-VidCRAFT3 accurately reproduces intended object trajectories and maintains specified camera movements, closely matching the ground truth.
+Additional Qualitative Results
+This section presents additional qualitative comparisons, highlighting VidCRAFT3’s capabilities in controlling camera motion and object motion.
 Camera Motion Control.
@@
 16 compares VidCRAFT3’s results with state-of-the-art methods (CameraCtrl, MotionCtrl, CamI2V) and ground-truth sequences, highlighting superior camera trajectory accuracy and visual coherence.
+These examples demonstrate VidCRAFT3’s superior capability in generating visually coherent and precise camera motion across diverse scenarios.
 Object Motion Control.
@@
 17 provides comparisons against Image Conductor and Motion-I2V, showcasing VidCRAFT3’s superior capability in generating realistic object trajectories and visually coherent motion across diverse scenarios.
-Lighting Direction Control.
-Fig.
-15 further illustrates VidCRAFT3’s precise control over lighting direction, achieving accurate shading, shadows, and reflections, consistent with ground-truth sequences.
+The examples demonstrate VidCRAFT3’s improved capability in accurately reproducing specified object trajectories, achieving more realistic and coherent object motion across diverse scenarios.
 These additional qualitative evaluations highlight VidCRAFT3’s comprehensive and precise control capabilities, confirming its robustness and versatility across different control dimensions.
-Appendix E Limitations
-VidCRAFT3 can generate unstable results for many specific situations, e.g., large human motion, phyisical interactions, significant change of lighting conditions.
-This is mostly due to the lack of diverse training data and we believe the neural architecture can also be improved to promote a better understanding of physics and 3D spatial relationships of objects.
-The camera motion control and object motion control datasets contain inaccurate annotations of camera pose and motion trajectories, which sometimes result in blurred generated videos.
-Currently, VidCRAFT3 only offers control over the lighting direction and therefore can be improved to support full HDR or other more fine-grained representations of the light field.
-Figure 8: Qualitative comparisons of training strategies (Dense, Sparse, Dense+Sparse) in the ablation study conducted on WebVid-10M.
-Results clearly illustrate that Dense training struggles with alignment given sparse input trajectories during inference, Sparse training improves alignment but loses motion detail, while Dense+Sparse training effectively combines both strategies, resulting in superior alignment and enhanced motion realism compared to the ground truth (GT).
-Figure 9: Qualitative comparison of lighting embedding integration strategies (Text Cross-Attn, Time Embedding, Lighting Cross-Attn) in the ablation study on the VLD dataset.
-Results clearly demonstrate that the Lighting Cross-Attn strategy provides more accurate and realistic lighting control, achieving superior shadows, reflections, and overall visual fidelity compared to Text Cross-Attn and Time Embedding, aligning closely with the ground truth (GT).
-Figure 10: Qualitative comparison between Fourier Embedding and SH Encoding for representing lighting direction on the VLD dataset.
-Results demonstrate that SH Encoding consistently produces more realistic lighting effects, achieving better shading accuracy and reflection details compared to Fourier Embedding, aligning closely with the ground truth (GT).
-Figure 11: Additional qualitative results demonstrating simultaneous control over camera motion and lighting direction.
-Each pair of rows compares the original video with the corresponding relighting results generated by VidCRAFT3, highlighting the framework’s effectiveness in precisely adjusting lighting direction while preserving accurate camera motion.
-Figure 12: Additional qualitative results showcasing simultaneous camera motion and object motion control.
-Each example compares our generated video (Ours) with ground-truth videos (GT), demonstrating VidCRAFT3’s capability to accurately reproduce intended object motion while precisely maintaining the specified camera motion.
+Limitations
+VidCRAFT3 can generate unstable results for many specific scenarios, e.g., large human motion, physical interactions, and significant changes of lighting conditions.
+This is mainly due to the lack of diverse training data.
+Additionally, we believe the neural architecture can be further improved to enhance the model’s understanding of physics and 3D spatial relationships.
+Inaccurate annotations in the camera motion control and object motion control datasets, particularly for camera poses and motion trajectories, can sometimes lead to blurred videos.
+Currently, VidCRAFT3 only offers control over lighting direction.
+To improve, the model could be extended to support full HDR or other more fine-grained representations of the light field.
+[…]
+Figure 14: Qualitative comparison of lighting embedding integration strategies (Text Cross-Attn, Time Embedding, and Lighting Cross-Attn) in the ablation study on the VLD dataset.
+Each example shows frames generated under different lighting embedding strategies.
+Text Cross-Attn integrates lighting embeddings with text embeddings through cross-attention but results in less accurate lighting control.
+Time Embedding improves consistency by incorporating lighting information into time embedding, but still lacks precision.
+Lighting Cross-Attn directly integrates lighting embeddings with dedicated cross-attention, providing superior control over lighting directions, producing more realistic shadows, reflections, and overall lighting effects, and aligning more closely with the ground truth (GT) compared to the other strategies.
+Figure 15: Qualitative comparison between Fourier Embedding and SH Encoding for representing lighting direction in the ablation study on the VLD dataset.
+Each example shows frames generated under different lighting direction representations.
+Fourier Embedding represents lighting direction using periodic basis functions but struggles with accurately capturing complex lighting effects.
+SH Encoding utilizes spherical harmonics to better model lighting direction, producing more realistic and detailed lighting effects, including enhanced shading and reflections.
+SH Encoding aligns more closely with the ground truth (GT), providing superior visual fidelity and more accurate lighting control compared to Fourier Embedding.
 Figure 16: Additional qualitative comparisons for camera motion control.
@@
 The examples demonstrate VidCRAFT3’s improved capability in accurately reproducing specified object trajectories, achieving more realistic and coherent object motion across diverse scenarios.
-Figure 15: Additional qualitative results demonstrating lighting direction control.
-Our generated videos (”Ours”) are compared against ground-truth sequences (”GT”), clearly highlighting VidCRAFT3’s capability to accurately reproduce different lighting direction, achieving realistic shading, shadowing, and reflection details across diverse scenes.